### v1

# Finite solvable groups of square-free Euler characteristic

Abstract.
A characterization is completed for finite groups which act regularly on maps with square-free Euler characteristic, associated with infinite families of square-free Euler characteristic regular maps presented. This is based on a classification of finite groups of which each Sylow subgroup has a cyclic or dihedral subgroup of prime index.
Keywords: solvable group, regular map, square-free, Euler characteristic
The project was partially supported by the NNSF of China (11931005).
November 18, 2025
1. Introduction
Edge-transitive maps are categorized into fourteen classes according to local structures and local actions in [9, 15]. The problem of constructing and classifying important classes of such maps with specific Euler characteristic has attracted considerable attention; refer to [4] - [7] for the Euler characteristic to be a negative prime or product of two primes. We aim towards to a classification of regular maps with square-free Euler characteristic.
Assume that $G$ is a finite group acting edge-transitively on a map with square-free Euler characteristic, which is said to be of square-free Euler characteristic for convenience. Then $G$ has the following restricted property (Lemma 2.1):
each Sylow subgroup has a cyclic or dihedral subgroup of prime index.
It leads us to studying groups of square-free Euler characteristic. This project extends the characterization of almost-Sylow cyclic groups, namely, each Sylow subgroup of odd order is cyclic, and the Sylow $2$-subgroup either is trivial or contains a cyclic subgroup of index 2; refer to Zassenhaus [18], Suzuki [16] and Wong [17].
In previous work [11], we characterized non-solvable groups of square-free Euler characteristic, and in [10], we determined $2$-groups of square-free Euler characteristic. As the third paper of this series, the purpose of this paper is to present a classification (characterization) of such groups in the solvable case.
As usual, for a group $G$ and a subset $\pi$ of prime divisors of $|G|$, denote by $G_{\pi}$ a Hall $\pi$-subgroup of $G$, which is a Sylow subgroup if $\pi$ contains a single prime $p$, and denoted by $G_{p}$. For a positive integer $n$, denote by ${\sf Z}_{n}$ a cyclic group of order $n$, by ${\rm D}_{2n}$ or ${\rm Q}_{2n}$ a dihedral or generalized quaternion group of order $2n$, and by ${\rm A}_{n}$ or ${\rm S}_{n}$ an alternating or symmetric group of degree $n$, respectively. A finite group $G$ is a semidirect product of $A$ by $B$ if $A\lhd G$ and $B<G$ are such that $G/A\cong B$, denoted by G=A:BG=A{:}B. Our first main result is the following theorem, which characterizes finite solvable groups of which each Sylow subgroup has a cyclic or dihedral subgroup of prime index.
Theorem 1.1.
Let $G$ be a finite solvable group such that each Sylow subgroup of $G$ has a cyclic or dihedral subgroup of prime index. Then:
(1) The Sylow subgroups of $G$ are characterized in Proposition 4.1 (for $G_{2}$) and Lemma 4.3 (for $G_{p}$, $p$ odd prime).
(2) G=(A:B):KG=(A{:}B){:}K, where $\gcd(|A{:}B|,|K|)=1$, such that
  • (i)
    AA is abelian, $B$ is nilpotent, $A\cap{\bf Z}(A{:}B)=1;$
  • (ii)
    $K$ is one of the following groups:
    • (a)
      $K=G_{2}$ ($K$ might be $1$ in case of $G$ of odd order);
    • (b)
      $K=G_{\{2,3\}}$, explicitly given in Lemma 5.6 (ii) - (iii);
    • (c)
      $K=G_{\{2,7\}}={\sf Z}_{2}^{3}{:}G_{7}$, or $K=G_{\{2,3,7\}}={\sf Z}_{2}^{3}{:}(G_{7}{:}G_{3})$.
The next theorem shows that automorphism groups of regular maps with square-free Euler characteristic are more restricted.
Theorem 1.2.
Let $\mathcal{M}$} be a regular map with square-free Euler characteristic, and let $G={\sf Aut}{\mathcal{M}}$ be solvable. Then G=(A:B):KG=(A{:}B){:}K, where $A$ is abelian, $B$ is nilpotent, $A\cap{\bf Z}(A{:}B)=1$, and $K=G_{2}$ or $G_{\{2,3\}}$ is one of the following groups:
  1. (i)
    $K={\sf Z}_{2}$, ${\rm D}_{2^{\ell+1}}$, ${\rm D}_{2^{\ell+1}}\times{\sf Z}_{2}$ or ${\rm D}_{2^{\ell+3}}{:}{\sf Z}_{2}$, where $\ell\geqslant 1;$
  2. (ii)
    $K=X$ or $X\times{\sf Z}_{2}$, where $X\in\{{\rm S}_{4},\,{\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}},\,({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2},\,({\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}\};$
  3. (iii)
    $K={\rm D}_{2.3^{\ell}}\times{\rm S}_{4}$ or $({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\rm D}_{6}$.
The paper is organized as follows. After the Introduction, we collect some preliminaries in Section 2, and construct several infinite families of regular maps with square-free Euler characteristic in Section 3. Then, we prove our two main results in Section 4 - 5 and Section 6, respectively.
2. Preliminaries
Let $G$ be a group. Denote by $\Phi(G)$ and $G^{\prime}$ the Frattini subgroup and the commutator subgroup of $G$. If $G$ is a $p$-group, then $^{\prime}\leqslant\Phi(G), G/G$$G^{\prime}$ is abelian, and $G/\Phi(G)$ is elementary abelian. For a subgroup $H\leqslant G$, denote by ${\bf N}_{G}(H)$ and ${\bf C}_{G}(H)$ the normalizer and the centralizer of $H$ in $G$, respectively.
Let $\it\Gamma$} be a graph and $\lambda$ be a positive integer. Denote by ${\it\Gamma}^{(\lambda)}$ the graph with each edge of $\it\Gamma$} replaced by $\lambda$ edges.
It is known (refer to [12]) that an arc-transitive (multi-) graph $\it\Gamma$} can be expressed as a coset graph :
${\it\Gamma}={\sf Cos}(G,H,J),\ \mbox{or}\ {\sf Cos}(G,H,HgH)\ \mbox{(for simple graphs)},$
which has vertex set $[G:H]$ and edge set $[G:J]$, such that
    $Hg_{1}$ is incident with $Jg_{0}$ if and only if $g_{1}g_{0}^{-1}\in HJ$; or equivalently,
    $Hg_{1}$, $Hg_{2}$ are adjacent if and only if $g_{2}g_{1}^{-1}\in HgH$ in the simple graph case.
A map $\mathcal{M}$} is a $2$-cell embedding of a graph into a closed surface. Let ${\mathcal{M}}=(V,E,F)$ be a map with vertex set $V$, edge set $E$ and face set $F$, and let $G={\sf Aut}{\mathcal{M}}$. The Euler characteristic of $\mathcal{M}$} is given by Euler formula:
$\chi({\mathcal{M}})=|V|-|E|+|F|.$
A flag of $\mathcal{M}$} is an incidence triple $(\alpha,e,f)$ with $\alpha\in V$, $e\in E$ and $f\in F$. Given a flag $(\alpha,e,f)$, it is well-known that each of the stabilizers $G_{\alpha}$ and $G_{f}$ is cyclic or dihedral, and meanwhile, $G_{e}=1$, ${\sf Z}_{2}$ or ${\rm D}_{4}$.
Further, if $\mathcal{M}$} is a flag-regular map (also called a regular map), i.e., $G$ acts regularly on flags, then there exists a triple $(x,y,z)$ of involutions of $G$ which determines the map $\mathcal{M}$} as follows:
$G_{\alpha}=\langle x,y\rangle$, $G_{e}=\langle x,z\rangle\cong{\rm D}_{4}$, $G_{f}=\langle y,z\rangle$.
Meanwhile, the Euler characteristic has the form
$\mbox{$\chi({\mathcal{M}})={|G|\over|G_{\alpha}|}-{|G|\over 4}+{|G|\over|G_{f}|}$}.$
Conversely, if a group $G$ has a triple of involutions $(x,y,z)$ such that
$G=\langle x,y,z\rangle$ and $xz=zx$ $(x\neq z)$,
then there exists a unique regular map ${\mathcal{M}}=(V,E,F)$ with
$V=[G:\langle x,y\rangle],\ E=[G:\langle x,z\rangle],\ F=[G:\langle y,z\rangle],$
and the incidence relation defined by non-empty intersection, such that $G={\sf Aut}{\mathcal{M}}$ acts regularly on flags. Hence, such a triple of involutions is called a regular triple for $G$, and the induced regular map is denoted by ${\mathcal{M}}(G,x,y,z)$. Moreover, the underlying graph of ${\mathcal{M}}(G,x,y,z)$ can be expressed as a coset graph:
${\sf Cos}(G,\langle x,y\rangle,\langle x,z\rangle)$, or ${\sf Cos}(G,H,HzH)$ where $H=\langle x,y\rangle$ (for simple graphs).
In conclusion, each of the following two things determines the other:
    the regular maps arisen from a group $G$;
    the regular triples for $G$.
The following lemma established in [11] characterizes Sylow subgroups of $G={\sf Aut}({\mathcal{M}})$ in the case where $\mathcal{M}$} is with square-free characteristic. We note that if $\chi({\mathcal{M}})$ is square-free, then of course so is $\gcd(\chi({\mathcal{M}}),|{\sf Aut}{\mathcal{M}}|)$.
Lemma 2.1.
([11, Lem. 2.2]) Let $\mathcal{M}$} be a map, and let $G={\sf Aut}{\mathcal{M}}$. Assume that $\gcd(\chi({\mathcal{M}}),|G|)$ is square-free. Then each Sylow $p$-subgroup of $G$ has a cyclic or dihedral subgroup of index $p$.
We thus study finite groups satisfying the following hypothesis.
Hypothesis 2.2.
Let $G$ be a finite group such that each Sylow subgroup of $G$ has a cyclic or dihedral subgroup of prime index.
The next lemma plays a key role in the analysis of groups satisfying Hypothesis 2.2.
Lemma 2.3.
([11, Lem. 2.5]) Let a finite group $G$ satisfy Hypothesis 2.2. Then so does each subgroup and each factor group of $G$.
3. Examples of square-free Euler characteristic regular maps
We construct several infinite families of regular maps with square-free Euler characteristic in this section. The constructions are based on the following groups:
$\begin{array}[]{lllll}X&=&(\langle a,s\rangle\times\langle b,t\rangle){:}\langle\sigma\rangle\cong{\rm D}_{2n}\wr{\rm S}_{2}&&\\ &=&\langle a,b\rangle{:}\langle s,\sigma\rangle\cong{\sf Z}_{n}^{2}{:}{\rm D}_{8},&&(\ast)\end{array}$
where $\langle a,s\rangle=\langle a\rangle{:}\langle s\rangle={\rm D}_{2n}$, $\langle b,t\rangle=\langle b\rangle{:}\langle t\rangle={\rm D}_{2n}$, $(a,s)^{\sigma}=(b,t)$ and $|\sigma|=2$.
The first family is given now.
Construction 3.1.
Let $G=\langle a,s\rangle\times\langle b,t\rangle={\rm D}_{2n}\times{\rm D}_{2n}\leqslant X$ as given in ($\ast$), with $n$ odd. Let
x=s,y=abst,z=st.x=s,\ y=abst,\ z=st.
Then $(x,y,z)$ is a triple of involutions with $xz=zx$. Further,
$(yx)^{2}=(abt)^{2}=a^{2}$ and $(yxz)^{2}=(abs)^{2}=b^{2}$.
As $n$ is odd, $\langle a\rangle=\langle a^{2}\rangle$ and $\langle b\rangle=\langle b^{2}\rangle$, so that $G=\langle x,y,z\rangle$. Thus, $(x,y,z)$ is a regular triple for $G$.
Lemma 3.2.
Let ${\mathcal{M}}_{n}={\mathcal{M}}(G,x,y,z)$ be a regular map defined in Construction 3.1. Then the underlying graph of ${\mathcal{M}}_{n}$ is a multiple cycle ${\bf C}_{n}^{(n)}$, and $\chi({\mathcal{M}}_{n})=-n(n-3)$. In particular, $\chi({\mathcal{M}}_{n})$ is square-free for infinitely many odd values of $n$.
Proof.
The vertex stabilizer $H$ is
$H=\langle x,y\rangle=\langle s,abst\rangle=\langle abt\rangle{:}\langle s\rangle=\langle a,s\rangle\times\langle bt\rangle={\rm D}_{4n}.$
Meanwhile, the face stabilizer is
$\langle y,z\rangle=\langle abst,st\rangle=\langle ab\rangle{:}\langle st\rangle={\rm D}_{2n}.$
Thus the map ${\mathcal{M}}_{n}$ has valency $2n$ and face length $n$. Further, $H\cap H^{z}=\langle a,s\rangle\cong{\rm D}_{2n}$ is of index $2$ in $H$, so the underlying graph ${\it\Gamma}={\sf Cos}(G,H,HzH)\cong{\bf C}_{n}^{(n)}$.
The Euler characteristic
$\chi({\mathcal{M}}_{n})=|G|\left({1\over|\langle x,y\rangle|}-{1\over 4}+{1\over|\langle y,z\rangle|}\right)=4n^{2}\left({1\over 4n}-{1\over 4}+{1\over 2n}\right)=-n(n-3).$
Let $n=4x+1$. Then $\chi({\mathcal{M}}_{n})=-2(4x+1)(2x-1)$. By a result given by G. Ricci in 1933 (which is improved by $\rm P.Erd\ddot{o}s$} in [8]), there are infinitely many positive integers $x$ for which f(x)=(4x+1)(2x1)f(x)=(4x+1)(2x-1) is square-free, so there are infinitely many odd values of $n$ such that $\chi({\mathcal{M}}_{n})$ is square-free. $\Box$
Here is some data such that n(n3)-n(n-3) is square-free, with $n$ odd.
$\begin{array}[]{c|ccccccc}\hline\cr n&5&13&17&29&37&41&\cdots\\ -n(n-3)&-2.5&-2.5.13&-2.7.17&-2.13.29&-2.17.37&-2.19.41$$\cdots\\ \hline\cr\end{array}$
Next we construct the second family.
Construction 3.3.
Let $G=\langle a,b\rangle{:}\langle st,\sigma\rangle={\sf Z}_{n}^{2}{:}{\sf Z}_{2}^{2}\leqslant X$ as given in ($\ast$). Let
$x=\sigma,\ y=ast,\ z=abst.$
Then $(x,y,z)$ is a triple of involutions with $xz=zx$. Further,
$yz=b^{-1}$ and $(yz)^{x}=a^{-1}$,
so that $G=\langle x,y,z\rangle$. Thus, $(x,y,z)$ is a regular triple for $G$.
Lemma 3.4.
Let ${\mathcal{M}}_{n}={\mathcal{M}}(G,x,y,z)$ be a regular map defined in Construction 3.3. Then the underlying graph of ${\mathcal{M}}_{n}$ is ${\bf C}_{n}^{(n)}$, and $\chi({\mathcal{M}}_{n})=-n(n-3)$. In particular, $\chi({\mathcal{M}}_{n})$ is square-free for infinitely many values of $n$.
Proof.
The vertex stabilizer $H$ is
$H=\langle x,y\rangle=\langle ast\sigma\rangle{:}\langle\sigma\rangle={\rm D}_{4n}.$
Meanwhile, the face stabilizer is
$\langle y,z\rangle=\langle abst,ast\rangle=\langle b\rangle{:}\langle ast\rangle={\rm D}_{2n}.$
Thus the map ${\mathcal{M}}_{n}$ has valency $2n$ and face length $n$. Further, $H\cap H^{z}=\langle a^{-1}b\rangle{:}\langle\sigma\rangle={\rm D}_{2n}$ is of index $2$ in $H$, so the underlying graph ${\it\Gamma}={\sf Cos}(G,H,HzH)\cong{\bf C}_{n}^{(n)}$.
The Euler characteristic
$\chi({\mathcal{M}})=|G|\left({1\over|\langle x,y\rangle|}-{1\over 4}+{1\over|\langle y,z\rangle|}\right)=4n^{2}\left({1\over 4n}-{1\over 4}+{1\over 2n}\right)=-n(n-3).$
As Lemma 3.2 shows, there are infinitely many odd values of $n$ such that $\chi({\mathcal{M}}_{n})$ is square-free. Further, let $n=4x+2$. Then $\chi({\mathcal{M}}_{n})=-2(2x+1)(4x-1)$. According to [8], there are infinitely many positive integers $x$ for which f(x)=(2x+1)(4x1)f(x)=(2x+1)(4x-1) is square-free, so there are infinitely many even values of $n$ such that $\chi({\mathcal{M}}_{n})$ is square-free. $\Box$
Here is some data such that n(n3)-n(n-3) is square-free, with $n$ even.
$\begin{array}[]{c|ccccccc}\hline\cr n&2&10&14&22&26&34&\cdots\\ -n(n-3)&2&-2.5.7&-2.7.11&-2.11.19&-2.13.23&-2.17.31$$\cdots\\ \hline\cr\end{array}$
The third family is given below.
Construction 3.5.
Let $G=\langle a,b\rangle{:}\langle s,\sigma\rangle={\sf Z}_{n}^{2}{:}{\rm D}_{8}=X$ as given in ($\ast$), with $n$ odd. Let
$x=\sigma,\ y=s,\ z=abst.$
Then $(x,y,z)$ is a triple of involutions with $xz=zx$. Further,
$(zy)^{2}=(abt)^{2}=a^{2}$ and $((zy)^{2})^{x}=(a^{2})^{x}=b^{2}$.
As $n$ is odd, $\langle a\rangle=\langle a^{2}\rangle$ and $\langle b\rangle=\langle b^{2}\rangle$, so that $G=\langle x,y,z\rangle$. Thus, $(x,y,z)$ is a regular triple for $G$.
Lemma 3.6.
Let ${\mathcal{M}}_{n}={\mathcal{M}}(G,x,y,z)$ be a regular map defined in Construction 3.5. Then the underlying graph of ${\mathcal{M}}_{n}$ is ${\bf C}_{n}\square{\bf C}_{n}$, and $\chi({\mathcal{M}}_{n})=-n(n-2)$. In particular, $\chi({\mathcal{M}}_{n})$ is square-free for infinitely many prime values of $n$.
Proof.
The vertex stabilizer $H$ is
$H=\langle x,y\rangle=\langle\sigma,s\rangle={\rm D}_{8}.$
Meanwhile, the face stabilizer
$\langle y,z\rangle=\langle s,abst\rangle=\langle abt\rangle{:}\langle s\rangle={\rm D}_{4n}.$
Thus the map ${\mathcal{M}}_{n}$ has valency $4$ and face length $2n$. Let ${\it\Gamma}={\sf Cos}(G,H,HzH)$ be the underlying graph of ${\mathcal{M}}_{n}$. Then $G=\langle a,b\rangle{:}H$, and so $\langle a,b\rangle$ is regular on the vertex set of $\it\Gamma$}. Hence $\it\Gamma$} is a (connected and undirected) Cayley graph ${\sf Cay}(K,S)$ of valency 4, where $K=\langle a,b\rangle={\sf Z}_{n}^{2}$. It follows that $S=\{x,x^{-1},y,y^{-1}\}$ with $\langle x,y\rangle=\langle a,b\rangle$, and so ${\it\Gamma}={\bf C}_{n}\square{\bf C}_{n}$.
The Euler characteristic
$\chi({\mathcal{M}}_{n})=|G|\left({1\over|\langle x,y\rangle|}-{1\over 4}+{1\over|\langle y,z\rangle|}\right)=8n^{2}\left({1\over 8}-{1\over 4}+{1\over 4n}\right)={-n(n-2).}$
According to [8], there are infinitely many odd values of $n$ such that $\chi({\mathcal{M}}_{n})$ is square-free. Further, by a theorem of Chen [3], there indeed exist infinitely many primes $p$ such that p2p-2 is square-free, then so is p(p2)-p(p-2). $\Box$
Here is some data such that p(p2)-p(p-2) is square-free, with $p$ prime.
$\begin{array}[]{c|cccccccc}\hline\cr p&3&5&7&13&17&19&23&\cdots\\ -p(p-2)&-3&-3.5&-5.7&-11.13&-3.5.17&-17.19&-3.7.23$$\cdots\\ \hline\cr\end{array}$
4. $p$-Groups
In this section, we study $p$-groups satisfying Hypothesis 2.2, i.e., $p$-groups with a cyclic or dihedral subgroup of index $p$. Note that the case for $p=2$ is dealt with in the previous work [10], which is introduced below.
Let $A,B$ be finite groups whose centers ${\bf Z}(A),{\bf Z}(B)$ have isomorphic subgroups $C_{1},C_{2}$, respectively. Let $\varphi$ be an isomorphism from $C_{1}$ to $C_{2}$, and let $C=\{(c,c^{\varphi})\mid c\in C_{1}\}$. Then $C$ is a normal subgroup of $A\times B$. The factor group $(A\times B)/C$ is called a central product of $A$ and $B$, denoted by $A\circ_{C}B$, and sometimes simply by $A\circ B$ if $C$ is equal to ${\bf Z}(A)$ or ${\bf Z}(B)$.
Let ${\rm Q}_{2^{\ell+1}}=\langle u,v\rangle$ be a generalized quaternion group of order $2^{\ell+1}$, where
$|u|=2^{\ell}\geqslant 2^{2}$, $u^{2^{\ell-1}}=v^{2}$, and $u^{v}=u^{-1}$.
Then ${\rm Q}_{2^{\ell+1}}$ has $u^{2^{\ell-1}}=v^{2}$ as the only involution. Let
$G={\rm Q}_{2^{\ell+1}}\circ{\sf Z}_{4}=\langle u,v\rangle\circ\langle b\rangle=(\langle u,v\rangle\times\langle b\rangle)/\langle b^{2}v^{2}\rangle.$
Then $G$ contains normal subgroups $\langle u,v\rangle={\rm Q}_{2^{\ell+1}}$ and $\langle b\rangle={\sf Z}_{4}$, and the only involutions contained in $\langle u,v\rangle$ and $\langle b\rangle$ are $u^{2^{\ell-1}}=v^{2}$ and $b^{2}$, respectively, which are actually identical in $G$ as $b^{2}v^{2}=1$.
Proposition 4.1.
([10], Thm. 1.1) Let $G$ be a $2$-group which has a cyclic or dihedral subgroup of index $2$. Then $G$ is one of the following groups. Further, either ${\sf Aut}(G)$ is a $2$-group, or $(G,{\sf Aut}(G))$ is one of $({\sf Z}_{2}^{2},{\rm S}_{3}),\,({\sf Z}_{2}^{3},{\rm GL}_{3}(2)),\,({\rm Q}_{8},{\rm S}_{4}),\,({\rm Q}_{8}\circ{\sf Z}_{4},{\rm S}_{4}\times{\sf Z}_{2})$.
  • (i)
    $G$ is one of the following groups, with a cyclic subgroup of index $2$, $\ell\geqslant 1$:
    • (i.1)
      $G={\sf Z}_{2^{\ell}}$ or ${\sf Z}_{2^{\ell}}\times{\sf Z}_{2};$
    • (i.2)
      $G={\rm D}_{2^{\ell+2}}$ or ${\rm Q}_{2^{\ell+2}};$
    • (i.3)
      $G={\sf Z}_{2^{\ell+2}}{:}{\sf Z}_{2}=\langle a\rangle{:}\langle b\rangle$ with $a^{b}=a^{2^{\ell+1}\pm 1};$
  • (ii)
    $G$ is one of the following groups, with a dihedral subgroup of index $2$, $\ell\geqslant 1$:
    • (ii.1)
      $G={\rm D}_{2^{\ell+1}}\times{\sf Z}_{2};$
    • (ii.2)
      $G={\rm D}_{2^{\ell+3}}{:}{\sf Z}_{2}=\langle a,b\rangle{:}\langle c\rangle$, where $\langle a\rangle{:}\langle b\rangle={\rm D}_{2^{\ell+3}}$, $(a,b)^{c}=(a^{2^{\ell+1}+1},b);$
    • (ii.3)
      $G={\rm Q}_{2^{\ell+2}}\circ{\sf Z}_{4}$.
Proposition 4.2.
([10], Cor. 4.9) Let $G$ be a $2$-group which has a cyclic or dihedral subgroup of index $2$. If $G$ has a regular triple, then it is one of the following groups:
  • (i)
    $G={\rm D}_{2^{\ell+1}}$ or ${\rm D}_{2^{\ell+1}}\times{\sf Z}_{2}$, $\ell\geqslant 1;$
  • (ii)
    $G={\rm D}_{2^{\ell+3}}{:}{\sf Z}_{2}=\langle a,b\rangle{:}\langle c\rangle$, where $\langle a\rangle{:}\langle b\rangle={\rm D}_{2^{\ell+3}}$, $(a,b)^{c}=(a^{2^{\ell+1}+1},b)$, $\ell\geqslant 1$.
For the odd primes $p$, we have the following characterization.
Lemma 4.3.
Let $p>2$ be an odd prime, and let $G$ be a $p$-group which has a cyclic subgroup of index $p$. Then one of the following holds. In particular, each prime divisor of $|{\sf Aut}(G)|$ is less than or equal to $p$.
  • (i)
    $G={\sf Z}_{p^{\ell}}$ is cyclic, and ${\sf Aut}(G)={\sf Z}_{(p-1)p^{\ell-1}}$.
  • (ii)
    $G={\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$ is abelian, and
    1. (ii.1)
      if $\ell>1$, then ${\sf Aut}(G)$ is solvable, and letting $G=\langle a,b\rangle$, then any Hall $p^{\prime}$-subgroup of ${\sf Aut}(G)$ is conjugate to $({\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle))_{p^{\prime}}={\sf Z}_{p-1}\times{\sf Z}_{p-1};$
    2. (ii.2)
      if $\ell=1$, then ${\sf Aut}(G)={\rm GL}(2,p)$, and further, any subgroup $M\leqslant{\rm GL}(2,p)$ of order coprime to $2p$ is an abelian group.
  • (iii)
    $G=\langle a\rangle{:}\langle b\rangle={\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}$ is non-abelian with $a^{b}=a^{p^{\ell-1}+1}$, $\ell\geqslant 2$, and ${\sf Aut}(G)$ is solvable such that any Hall $p^{\prime}$-subgroups of ${\sf Aut}(G)$ is conjugate to $\langle\rho\rangle={\sf Z}_{p-1}$, where $a^{\rho}=a^{\xi}\not=a$ with $\xi\,\,\big|\,\,p-1$, and $b^{\rho}=b$.
Proof.
According to [13, 5.3.4, pp.136], $G$ is one of the groups stated in this lemma. We only need to verify the statements regarding the automorphism groups of $G$.
(1) The automorphism groups of cyclic groups are well-known; refer to [14, Thm. 7.3].
(2) Suppose $G={\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$ with $\ell>1$. Then by [1, Thm. 3.2], ${\sf Aut}(G)$ is of order
$\begin{array}[]{ccl}|{\sf Aut}(G)|&=&|{\sf Aut}({\sf Z}_{p^{\ell}})|\cdot|{\sf Aut}({\sf Z}_{p})|\cdot|{\sf Hom}({\sf Z}_{p^{\ell}},{\sf Z}_{p})|\cdot|{\sf Hom}({\sf Z}_{p},{\sf Z}_{p^{\ell}})|\\ &=&(p-1)p^{\ell-1}\cdot(p-1)\cdot p\cdot p=p^{\ell+1}(p-1)^{2}.\end{array}$
More explicitly, let $G=\langle a\rangle\times\langle b\rangle$ with $|a|=p^{\ell}$, $|b|=p$. Then $\Phi(G)=\langle a^{p}\rangle$. Let $\overline{G}=G/\Phi(G)$, and consider the homomorphism
$\sigma:{\sf Aut}(G)\rightarrow{\sf Aut}(\overline{G})={\rm GL}(2,p).$
On the one hand, the kernel of $\sigma$ consists of automorphisms $g\in{\sf Aut}(G)$ which stabilizes each of the $\Phi(G)$-cosets, and in particular, stabilizes $\Phi(G)\langle a\rangle$ and $\Phi(G)\langle b\rangle$. Thus, by direct calculation we have:
$a^{g}\in\{a,a^{p+1},a^{2p+1},\cdots,a^{(p^{\ell-1}-1)p+1}\}\mbox{\ and\ }b^{g}\in\{b,a^{p^{\ell-1}}b,a^{2p^{\ell-1}}b,\cdots,a^{(p-1)p^{\ell-1}}b\}.$
Consequently, ${\rm Ker}(\sigma)=\langle x\rangle\times\langle y\rangle\cong{\sf Z}_{p^{\ell-1}}\times{\sf Z}_{p},$ where
$x:a\rightarrow a^{p+1},b\rightarrow b\ \mbox{\,and\,}\ y:a\rightarrow a,b\rightarrow a^{p^{\ell-1}}b.$
On the other hand, the image of $\sigma$ falls into the (solvable) parabolic subgroup $P$ of ${\sf Aut}(\overline{G})={\rm GL}(2,p)$ which stabilizes $\langle\overline{b}\rangle$, because
$\langle\overline{b}\rangle=\{\overline{1},\overline{b},\cdots,\overline{b}^{p-1}\}=\{\Phi(G),b\Phi(G),\cdots,b^{p-1}\Phi(G)\}$
exactly consists of those elements of $G$ of order less than $p^{\ell}$. Moreover, it follows from ${\sf Aut}(G)/{\rm Ker}(\sigma)\cong{\rm Im}(\sigma)$ that $|{\rm Im}(\sigma)|=p(p-1)^{2}=|P|$, and so ${\rm Im}(\sigma)=P$.
Then ${\sf Aut}(G)$ is solvable, since both ${\rm Ker}(\sigma)$ and ${\rm Im}(\sigma)$ are solvable. Note that the following subgroup
$X=({\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle))_{p^{\prime}}={\sf Z}_{p-1}\times{\sf Z}_{p-1}$
is a Hall $p^{\prime}$-subgroup of $G$. Then any other Hall $p^{\prime}$-subgroup is conjugate to $X$.
(3) Suppose $G={\sf Z}_{p}\times{\sf Z}_{p}$. Then ${\sf Aut}(G)={\rm GL}(2,p)$, of order $p(p+1)(p-1)^{2}$ with prime divisors less than or equal to $p$. Further, let $M\leqslant{\rm GL}(2,p)$ be a subgroup of order coprime to $2p$, and let $N=M/{\bf Z}(M)$. Then
$N=M/{\bf Z}(M)\leqslant M/({\bf Z}({\rm GL}(2,p))\cap M)\lesssim{\rm PGL}(2,p).$
Further, since ${\rm PSL}(2,p)N\leqslant{\rm PGL}(2,p)$, we have
$|N/({\rm PSL}(2,p)\cap N)|\leqslant|{\rm PGL}(2,p)|/|{\rm PSL}(2,p)|=2.$
It then follows by $2\nmid|N|$ that $N$ is indeed embedded into ${\rm PSL}(2,p)$. Note that the subgroups of ${\rm PSL}(2,p)$ are all explicitly known, by check which a subgroup of order coprime to $2p$ must be a cyclic group. Then $M$ is an abelian group, since $N=M/{\bf Z}(M)$ is cyclic.
(4) Suppose that $G=\langle a\rangle{:}\langle b\rangle={\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}$ with $a^{b}=a^{p^{\ell-1}+1}$ is non-abelian, $\ell\geqslant 2$. The automorphism group ${\sf Aut}(G)$ is described in [2, Thm. 3.1 & 3.2], by which we have: $|{\sf Aut}(G)|=(p-1)p^{\ell+1}$, ${\sf Aut}(G)$ is solvable with the Sylow $p$-subgroup normal, and $X=\langle\rho\rangle={\sf Z}_{p-1}$ is a Hall $p^{\prime}$-subgroup of $G$, where $\rho$ stabilizes $\langle a\rangle$ and centralizes $\langle b\rangle$, that is, $a^{\rho}=a^{\xi}\not=a$ with $\xi\,\,\big|\,\,p-1$, and $b^{\rho}=b$. $\Box$
5. Proof of Theorem 1.1
In this section, we study the way how the $p$-groups described in Section 4 compose a larger solvable group, and give a proof of Theorem 1.1
Let $G$ be a finite group, with $\{p_{1},\ldots,p_{s}\}$ the prime divisors of $|G|$. For $p=p_{i}$, denote by ${\bf O}_{p}(G)$ the maximal normal $p$-subgroup of $G$, or equivalently, the intersection of all Sylow $p$-subgroups of $G$. Let $F={\rm Fit}(G)$ be the Fitting subgroup of $G$, which is the largest nilpotent normal subgroup. Then
$F=F_{p_{1}}\times\cdots\times F_{p_{s}},\mbox{\ with\ each\ }F_{p_{i}}={\bf O}_{p_{i}}(G).$
Note that both ${\bf O}_{p}(G)$ and $F$ are characteristic in $G$. Further, if $G$ is solvable, then ${\bf C}_{G}(F)\leqslant F$, and so
$G/F={\bf N}_{G}(F)/\left(F{\bf C}_{G}(F)\right)\lesssim{\sf Out}(F).$
We first consider solvable groups of odd order.
Lemma 5.1.
Let $G$ be a solvable group of odd order satisfying Hypothesis 2.2. Let $p$ be the largest prime divisor of $|G|$. Then a Sylow $p$-subgroup $G_{p}$ is normal in $G$, and so $G=G_{p}{:}G_{p^{\prime}}$.
Proof.
Let $\overline{G}=G/{\bf O}_{p}(G)$, and let $F={\rm Fit}(\overline{G})$. Notice that ${\bf O}_{p}(\overline{G})=1$. We have $F=F_{p_{1}}\times\cdots\times F_{p_{s}}$ with all $p_{i}<p$. Now,
$\overline{G}/F\lesssim{\sf Out}(F)={\sf Out}(F_{p_{1}})\times\cdots\times{\sf Out}(F_{p_{s}}),$
and so $p\nmid|\overline{G}/F|$ by Lemma 4.3. Thus, $p\nmid|\overline{G}|$, so that, $G_{p}={\bf O}_{p}(G)\lhd G$. $\Box$
Lemma 5.2.
Let $G$ be a $\{p,q\}$-group satisfying Hypotheses 2.2, where $p>q$ are both odd primes, and $q\nmid p^{2}-1$. Then $G=G_{p}\times G_{q}$.
Proof.
It is proved in a similar way as the previous lemma. Note that by Lemma 4.3 any prime divisor of $|{\sf Aut}(G_{p})|$ divides p(p1)(p+1)p(p-1)(p+1). $\Box$
Before giving a general characterization of solvable groups of odd order, we need the following technical lemma.
Lemma 5.3.
Let $H={\sf Z}_{p^{\ell}}$ be a cyclic group and $x$ be an automorphism of $H$ with order coprime to $p$. Let $X=H{:}\langle x\rangle$. Then, (i) $H\leqslant X^{\prime};$ (ii) $H\cap{\bf Z}(X)=1$.
Proof.
By Lemma 4.3 (i), the order of $x$ divides p1p-1. Let $H=\langle a\rangle$ and suppose
$a^{x}=a^{m}$, where $1<m\leqslant p^{\ell}-1$,  $p\nmid m$  and  $p^{\ell}\mid m^{p-1}-1$.
If $p\mid m-1$, then by LTE (Lift the Exponent Lemma) we have $p^{\ell}\mid m-1$, a contradiction. Hence $p\nmid m-1$. Thus, (i) the commutator $a,x]=a^{-1}a^{x}=a^{m-1}$ can generate $H$, so that, $H\leqslant X^{\prime}$; (ii) for any $1\leqslant k<p^{\ell}$, $(a^{k})^{x}=a^{km}\neq a^{k}$, so that, $H\cap{\bf Z}(X)=1$. $\Box$
Lemma 5.4.
Let $G$ be a solvable group of odd order satisfying Hypothesis 2.2. Then G=C:DG=C{:}D, where $C$ is abelian, $D$ is nilpotent, and $C\cap{\bf Z}(G)=1$.
Proof.
Let $n$ be the number of prime divisors of $|G|$. The proof proceeds by induction on $n$. For $n=1$, the lemma holds with $C=1$, $D=G$. Suppose $n\geqslant 2$, and assume inductively that the lemma holds for groups of order with less prime divisors.
Let $p$ be the largest prime divisor of $|G|$. By Lemma 5.1 and the induction assumption, we have
$G=G_{p}{:}G_{p^{\prime}}=G_{p}{:}(A{:}B),$
with $A$ abelian, $B$ nilpotent, and $A\cap{\bf Z}(A{:}B)=1$, so that, $B$ does not centralize any Sylow subgroup of $A$. Further, if $G=G_{p}\times G_{p^{\prime}}$, then
$G=G_{p}{\times}(A{:}B)=(G_{p}{\times}A){:}B=A{:}(G_{p}\times B),$
which gives an expected factorization of $G$. We thus assume that $G\not=G_{p}\times G_{p^{\prime}}$, namely, $G_{p^{\prime}}$ acts by conjugation on $G_{p}$ non-trivially. We shall confirm
Claim: $G_{p}$ is centralized by each Sylow $r$-subgroup $A_{r}$ of $A$, so that, $G_{p}$ is centralized by $A$.
By Lemma 4.3, we have two cases to treat: $A_{r}={\sf Z}_{r^{k}}$ or ${\sf Z}_{r^{k}}\times{\sf Z}_{r}$.
Case 1. Let $A_{r}={\sf Z}_{r^{k}}$. If $A_{r},B_{r^{\prime}}]=1$, then by Lemma 4.3, $(A{:}B)_{r}=A_{r}{:}B_{r}={\sf Z}_{r^{k}}{:}{\sf Z}_{r}$ is non-abelian, and
$1\not=A_{r}\cap{\bf Z}(A_{r}{:}B_{r})=A_{r}\cap{\bf Z}(A{:}B)\leqslant A\cap{\bf Z}(A{:}B),$
a contradiction. Thus there exists a Sylow $q$-subgroup $B_{q}$ with $q\not=r$ which does not centralize $A_{r}$. Take $x\in B_{q}$ to be an element acting non-trivially on $A_{r}$ and let $X=A_{r}{:}\langle x\rangle$. Then $X$ acts on $G_{p}$ and induces an abelian automorphism group of $G_{p}$ by Lemma 4.3. It then follows by Lemma 5.3 (i) that $A_{r}\leqslant X^{\prime}\leqslant{\bf C}_{X}(G_{p})$.
Case 2. Let $A_{r}={\sf Z}_{r^{k}}\times{\sf Z}_{r}$, where $k\geqslant 1$. Then $r$ does not divide $|B|$. By Lemma 4.3 (ii), it implies that either
  • (i)
    $A_{r}={\sf Z}_{r}^{2}$ and $B$ is irreducible on $A_{r}$, or
  • (ii)
    there exists $a,b\in A_{r}$ such that $A_{r}=\langle a\rangle\times\langle b\rangle$, and both $\langle a\rangle$ and $\langle b\rangle$ are normalized by $B$.
For the former, $A_{r}$ centralizes $G_{p}$ since $A_{r}{:}B$ induces an abelian automorphism group of $G_{p}$ by Lemma 4.3. For the latter, there exist elements $u,v\in B$ such that $u$ acts on $\langle a\rangle$ non-trivially, and $v$ acts on $\langle b\rangle$ non-trivially. Then similar to Case 1 we have $G_{p},a]=[G_{p},b]=1$, so that, $G_{p},A_{r}]=[G_{p},\langle a,b\rangle]=1$.
We have therefore confirmed the claim. Consequently we have
$G=G_{p}{:}(A{:}B)=(G_{p}\times A){:}B.$
Meanwhile, by Lemma 4.3, $G_{p}$ is one of the following groups:
$G_{p}={\sf Z}_{p^{\ell}}$,  ${\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}$  or  ${\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$.
Note that by assumption, $G_{p^{\prime}}=A{:}B$ acts by conjugation on $G_{p}$ non-trivially, which implies that $B$ acts on $G_{p}$ non-trivially.
Suppose $G_{p}={\sf Z}_{p^{\ell}}$. As $B$ acts on $G_{p}$ non-trivially, it follows from Lemma 5.3 (ii) that $G_{p}\cap{\bf Z}(G)=1$. Thus G=C:DG=C{:}D satisfies the theorem, with $C=G_{p}\times A$, $D=B$.
Suppose $G_{p}={\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}=\langle a\rangle{:}\langle b\rangle$. By Lemma 4.3 (iii), $B$ acts on $\langle a\rangle$ non-trivially and acts on $\langle b\rangle$ trivially. Rewrite $G$ as
$G=(A\times(\langle a\rangle{:}\langle b\rangle)){:}B=(A\times\langle a\rangle){:}(B\times\langle b\rangle).$
Then by Lemma 5.3 (ii), G=C:DG=C{:}D satisfies the theorem, with $C=A\times\langle a\rangle$, $D=B\times\langle b\rangle$.
Suppose $G_{p}={\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$, where $\ell\geqslant 1$. By Lemma 4.3 (ii), it implies that either (i) $G_{p}={\sf Z}_{p}^{2}$ and $B$ is irreducible on $G_{p}$; or (ii) there exists $a,b\in G_{p}$ such that $G_{p}=\langle a\rangle\times\langle b\rangle$, and both $\langle a\rangle$ and $\langle b\rangle$ are normalized by $B$. For the former, G=C:DG=C{:}D satisfies the theorem, with $C=G_{p}\times A$, $D=B$. For the former, if $B$ acts non-trivially on both $\langle a\rangle$ and $\langle b\rangle$, then also let $C=G_{p}\times A$, $D=B$. Otherwise, $B$ acting non-trivially on $G_{p}$ implies $B$ acting non-trivially on some one of $\langle a\rangle$ and $\langle b\rangle$, say $\langle a\rangle$. Then similar to the last paragraph, let $C=A\times\langle a\rangle$, $D=B\times\langle b\rangle$. $\Box$
We next deal with solvable groups of even order.
Lemma 5.5.
Let $G$ be a solvable group satisfying Hypothesis 2.2. If $F:={\rm Fit}(G)={\bf O}_{2}(G)$, then either $G$ is a $2$-group, or $G$ is one of the following:
  • (i)
    $G={\sf Z}_{2}^{3}{:}{\sf Z}_{7}$ or ${\sf Z}_{2}^{3}{:}{\sf Z}_{7}{:}{\sf Z}_{3};$
  • (ii)
    $G={\sf Z}_{2}^{2}{:}H$, ${\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}H)$, ${\rm Q}_{8}{:}H$, or ${\sf Z}_{4}\circ({\rm Q}_{8}{:}H)$, where $H\in\{{\sf Z}_{3},{\rm S}_{3}\}$, ${\sf Z}_{2}^{2}{:}H\cong{\rm A}_{4}$ or ${\rm S}_{4}$, and ${\rm Q}_{8}{:}H\cong{\rm SL}(2,3)$ or ${\rm GL}(2,3)$.
Proof.
Assume that $G$ is not a 2-group. Then $G/F\leqslant{\sf Out}(F)$ is not a 2-group either. Thus by Lemma 4.1, $F$ is one of the following 2-groups:
$F\in\{{\sf Z}_{2}^{2},{\sf Z}_{2}^{3},{\rm Q}_{8},{\sf Z}_{4}\circ{\rm Q}_{8}\}.$
Moreover, ${\bf O}_{2}(G/F)$ is trivial, namely, $G/F$ has no normal 2-subgroups. Then by checking ${\sf Out}(F)$, either $|G/F|$ is divisible by 7, and $F={\sf Z}_{2}^{3}$, $G/F={\sf Z}_{7}$ or ${\sf Z}_{7}{:}{\sf Z}_{3}$; or $G/F={\sf Z}_{3}$ or ${\rm S}_{3}$. It follows that $G$ is one of the groups listed in the lemma. In particular, if $G=F{:}{\rm S}_{3}$, then $G_{2}=F{:}{\sf Z}_{2}>F$, and respectively for $F={\sf Z}_{2}^{2}$, ${\sf Z}_{2}^{3}$, ${\rm Q}_{8}$ or ${\sf Z}_{4}\circ{\rm Q}_{8}$, it is easy to verify that $G_{2}$ has a subgroup $K$ of index $2$ with $K={\sf Z}_{4}$, ${\rm D}_{8}$, ${\sf Z}_{8}$ or ${\rm D}_{16}$. That is, $G=F{:}{\rm S}_{3}$ surely satisfies Hypothesis 2.2. $\Box$
Lemma 5.6.
Let $K$ be a $\{2,p\}$-group with $p$ odd prime satisfying Hypotheses 2.2. If $K\neq K_{p}{:}K_{2}$, then one of the following holds.
  1. (i)
    $p=7$, $K={\sf Z}_{2}^{3}{:}K_{7};$
  2. (ii)
    $p=3$, $K=X{:}K_{3}$ with $X\in\{{\sf Z}_{2}^{2},\,{\sf Z}_{2}^{3},\,{\rm Q}_{8},\,{\sf Z}_{4}\circ{\rm Q}_{8}\};$
  3. (iii)
    $p=3$, $K$ is one of the groups in Tables 1 - 2
Proof.
Suppose that $K_{2}$ is normal in $K$, that is, $K=K_{2}{:}K_{p}$. Then $K_{2}$ has an automorphism of order $p$. By Lemma 4.1, $K_{2}$ is one of the following groups:
${\sf Z}_{2}^{2},\ {\rm Q}_{8},\ {\sf Z}_{2}^{3}\ \mbox{and}\ {\sf Z}_{4}\circ{\rm Q}_{8},$
and either $p=3$, $K=K_{2}{:}K_{3}$; or $K_{2}={\sf Z}_{2}^{3}$, $p=7$ and $K={\sf Z}_{2}^{3}{:}K_{7}$, as in part (i).
We then assume that neither $K_{p}$ nor $K_{2}$ is normal in $K$. Let
$F={\rm Fit}(K)=F_{p}\times F_{2}$, with $F_{p}<K_{p}$, $F_{2}<K_{2}$.
Consider the quotient group $\overline{K}=K/F_{p}$, whose Fitting group is a 2-group, namely ${\rm Fit}(\overline{K})={\bf O}_{2}(\overline{K})$. Thus, by Lemma 5.5, $\overline{K}={\bf O}_{2}(\overline{K}){:}H$, where either
(1) ${\bf O}_{2}(\overline{K})\in\{{\sf Z}_{2}^{2},{\sf Z}_{2}^{3},{\rm Q}_{8},{\sf Z}_{4}\circ{\rm Q}_{8}\}$, $H\in\{{\sf Z}_{3},{\rm S}_{3}\}$; or (2) ${\bf O}_{2}(\overline{K})={\sf Z}_{2}^{3}$, $H={\sf Z}_{7}$ or ${\sf Z}_{3}$.
Further, if (2) holds, $\overline{K}={\sf Z}_{2}^{3}{:}{\sf Z}_{p}$ with $p=3$ or $7$, then ${\sf Z}_{2}^{3}$ is a minimal normal subgroup of $\overline{K}$, and so $\overline{F_{2}}=F_{2}$ as a normal subgroup of $\overline{K}$ must be identity. Consequently, we have $F=F_{p}$, ${\sf Z}_{2}^{3}{:}{\sf Z}_{p}=K/F_{p}\lesssim{\sf Out}(F_{p})$, which does not hold for the $p$-groups appearing in Lemma 4.3.
Therefore, we have that (1) holds, and since $\overline{K}={\bf O}_{2}(\overline{K}){:}H$ is known, $F_{2}=\overline{F_{2}}$ as a normal 2-subgroup of $\overline{K}$ is also known. We then analyze all the possibilities.
Case 1. $F_{2}={\bf O}_{2}(\overline{K})$. In this case, we have $H={\rm S}_{3}$ since $F_{2}<K_{2}$. Then
$K=F_{3}.(F_{2}{:}{\rm S}_{3})=(F_{3}\times F_{2}).{\rm S}_{3}=F_{2}{:}(F_{3}.{\rm S}_{3})=F_{2}{:}(K_{3}{:}{\sf Z}_{2}).$
Note that $K_{3}{:}{\sf Z}_{2}/F_{3}\cong{\rm S}_{3}$. By checking the $3$-groups appearing in Lemma 4.3, we conclude that there are seven possibilities for the triple $T:=(K_{3},\,K_{3}{:}{\sf Z}_{2},\,F_{3})$ as follows (also note that ${\sf Aut}({\sf Z}_{3}\times{\sf Z}_{3})={\rm GL}_{2}(3)$ is solvable):
[…]
Consequently, Case 1 gives the groups listed in Table 1.
Table 1. Case: $F_{2}={\bf O}_{2}(\overline{K})$
case $F_{2}$ ${\sf Z}_{2}^{2}$ ${\sf Z}_{2}^{3}$ ${\rm Q}_{8}$ ${\sf Z}_{4}\circ{\rm Q}_{8}$
(1.1) ${\rm S}_{4}$ ${\sf Z}_{2}\times{\rm S}_{4}$ ${\rm GL}_{2}(3)$ ${\sf Z}_{4}\circ{\rm GL}_{2}(3)$
(1.2) ${\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}}$ ${\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}})$ ${\rm Q}_{8}{:}{\rm D}_{2.3^{\ell+1}}$ ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell+1}})$
(1.3) ${\sf Z}_{3^{\ell}}\times{\rm S}_{4},$ ${\sf Z}_{2.3^{\ell}}\times{\rm S}_{4}$ ${\sf Z}_{3^{\ell}}\times({\rm Q}_{8}{:}{\rm S}_{3})$ ${\sf Z}_{4.3^{\ell}}\circ({\rm Q}_{8}{:}{\rm S}_{3})$
(1.4) ${\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})$, ${\sf Z}_{6}\times({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})$ ${\sf Z}_{3}\times({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})$ ${\sf Z}_{12}\circ({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})$
(1.5) $({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2}$ $({\sf Z}_{2.3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2}$ $({\sf Z}_{3^{\ell}}\times({\rm Q}_{8}{:}{\rm S}_{3})){:}{\sf Z}_{2}$ $({\sf Z}_{4.3^{\ell}}\circ({\rm Q}_{8}{:}{\rm S}_{3})){:}{\sf Z}_{2}$
(1.6) $({\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}$ $({\sf Z}_{6}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}$ $({\sf Z}_{3}\times({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}$ $({\sf Z}_{12}\circ({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}$
(1.7) ${\sf Z}_{2}^{2}{:}({\rm D}_{2.3^{\ell}}{:}{\sf Z}_{3})$ ${\sf Z}_{2}\times{\sf Z}_{2}^{2}{:}({\rm D}_{2.3^{\ell}}{:}{\sf Z}_{3})$ ${\rm Q}_{8}{:}({\rm D}_{2.3^{\ell}}{:}{\sf Z}_{3})$ ${\sf Z}_{4}\circ({\rm Q}_{8}{:}({\rm D}_{2.3^{\ell}}{:}{\sf Z}_{3}))$
Case 2. $(F_{2},{\bf O}_{2}(\overline{K}))=({\sf Z}_{2}^{2},{\sf Z}_{2}^{3})$ or $({\rm Q}_{8},{\sf Z}_{4}\circ{\rm Q}_{8})$. Then
$K=F_{3}.({\bf O}_{2}(\overline{K}){:}H)=(F_{3}\times F_{2}).({\sf Z}_{2}\times H)=F_{2}{:}(F_{3}.({\sf Z}_{2}\times H)).$
Further, if $H={\sf Z}_{3}$, then $K/F_{2}=K_{3}{:}{\sf Z}_{2}$. Note that ${\sf Z}_{2}$ acts non-trivially on $K_{3}$, and meanwhile it acts trivially on $K_{3}/F_{3}=H$. Then by checking the $3$-groups appearing in Lemma 4.3, we conclude that there are three possibilities:
[…]
Otherwise, $H={\rm S}_{3}$, and so $K/F_{2}=K_{3}{:}({\sf Z}_{2}\times{\sf Z}_{2})$. By Lemma 4.3, ${\sf Z}_{2}\times{\sf Z}_{2}\leqslant{\sf Aut}(K_{3})$ implies $K_{3}={\sf Z}_{3^{\ell}}\times{\sf Z}_{3}$ and $K_{3}{:}({\sf Z}_{2}\times{\sf Z}_{2})={\rm D}_{2.3^{\ell}}\times{\rm D}_{6}$. We then get two possibilities:
[…]
Consequently, Case 2 gives the groups listed in Table 2.
Table 2. Case: $F_{2}<{\bf O}_{2}(\overline{K})$
case $(F_{2},{\bf O}_{2}(\overline{K}))$ $({\sf Z}_{2}^{2},{\sf Z}_{2}^{3})$ $({\rm Q}_{8},{\sf Z}_{4}\circ{\rm Q}_{8})$
(2.1) ${\rm D}_{2.3^{\ell}}\times{\rm A}_{4}$ $({\sf Z}_{3^{\ell}}{:}{\sf Z}_{4})\circ({\rm Q}_{8}{:}{\sf Z}_{3})$
(2.2) $({\sf Z}_{2}^{2}{:}{{\sf Z}_{3^{\ell}}})\times{\rm D}_{6}$ $({\rm Q}_{8}{:}{{\sf Z}_{3^{\ell}}})\circ({\sf Z}_{3}{:}{\sf Z}_{4})$
(2.3) $({\rm D}_{2.3^{\ell}}\times{\sf Z}_{2}^{2}){:}{\sf Z}_{3}$ $(({\sf Z}_{3^{\ell}}{:}{\sf Z}_{4})\circ{\rm Q}_{8}){:}{\sf Z}_{3}$
(2.4) ${\rm D}_{2.3^{\ell}}\times{\rm S}_{4}$ $({\sf Z}_{3^{\ell}}{:}{\sf Z}_{4})\circ({\rm Q}_{8}{:}{\rm S}_{3})$
(2.5) $({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\rm D}_{6}$ $({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})\circ({\sf Z}_{3}{:}{\sf Z}_{4})$
Case 3. Recall that $F_{2}$ is a normal 2-subgroup of ${\bf O}_{2}(\overline{K})$. Notice that ${\rm Q}_{8}{:}{\sf Z}_{3}$ or ${\rm Q}_{8}{:}{\rm S}_{3}$ only has the center ${\sf Z}_{2}$ and ${\rm Q}_{8}$ as normal $2$-subgroups. Then all the remaining cases are collected as follows :
[…]
For case (3.1), $(F_{2},{\bf O}_{2}(\overline{K}))=(1,{\rm Q}_{8})$, we have
$F=F_{3}\mbox{\ and\ }K/F_{3}={\rm Q}_{8}{:}H\lesssim{\sf Aut}(F_{3}).$
By lemma 4.3, it implies that $F_{3}={\sf Z}_{3}\times{\sf Z}_{3}$ and $K_{3}={\sf Z}_{9}{:}{\sf Z}_{3}$ is non-abelian. Let
$K_{3}=\langle a\rangle{:}\langle b\rangle={\sf Z}_{9}{:}{\sf Z}_{3}$, with $a^{b}=a^{4}$, $b^{a}=a^{6}b$,
where $F_{3}=\langle a^{3}\rangle\times\langle b\rangle$. Note that $K_{2}=({\rm Q}_{8}{:}H)_{2}={\rm Q}_{8}$ or ${\rm Q}_{8}{:}{\sf Z}_{2}$. Consider the central involution $c$ of $K_{2}$. Then $\overline{c}$ is the central involution of $\overline{K}$. On the one hand, $\overline{c}$ acts on $F_{3}$ by mapping each $x$ to $x^{-1}$. That is,
$x^{c}=x^{-1}$, so that, $c^{x}=x^{-2}c$ for any $x\in F_{3}$.
On the other hand, $\overline{c}^{\overline{a}}=\overline{c}$, and so $c^{a}=cy$ where $y\in F_{3}$. Let $y=a^{3i}b^{j}$. Then
$c^{a^{3}}=(ca^{3i}b^{j})^{a^{2}}=(ca^{(6i+6j)}b^{2j})^{a}=ca^{(9i+18j)}b^{3j}=c,\mbox{\ where\ }a^{3}\in F_{3},$
a contradiction. Hence case (3.1) cannot occur.
For case (3.2), $(F_{2},{\bf O}_{2}(\overline{K}))=(1,{\sf Z}_{4}\circ{\rm Q}_{8})$, we have ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\sf Z}_{3})\lesssim{\sf Aut}(F_{3})$, which is impossible by checking the 3-groups appearing in Lemma 4.3.
For case (3.3), since ${\bf O}_{2}(\overline{K})$ acts on $F_{3}$ with the centralizer ${\bf C}_{{\bf O}_{2}(\overline{K})}(F_{3})=F_{2}$, we deduce that in this case ${\sf Aut}(F_{3})$ contains a subgroup isomorphic to ${\rm A}_{4}\cong{\sf Z}_{2}^{2}{:}{\sf Z}_{3}$. By Lemma 4.3, $|{\sf Aut}(F_{3})|$ being divisible by $4$ implies $F_{3}={\sf Z}_{3^{\ell}}\times{\sf Z}_{3}$. For $\ell=1$, it is clear that ${\sf Aut}({\sf Z}_{3}^{2})={\rm GL}(3,2)\cong{\rm Q}_{8}{:}{\rm S}_{3}$ contains no subgroups isomorphic to ${\rm A}_{4}$. We thus assume $\ell>1$ and let $F_{3}=\langle a\rangle\times\langle b\rangle$ with $|a|=3^{\ell}$, $|b|=3$. By Lemma 4.3, ${\sf Aut}(F_{3})$ is solvable, and it contains a Hall $3^{\prime}$-subgroup
$H_{0}=({\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle))_{3^{\prime}}\cong{\sf Z}_{2}^{2},$
which stabilizes subgroups $\langle a\rangle$ and $\langle b\rangle$. Meanwhile, by simple calculation, $H_{0}$ cannot stabilize
    any of the other two subgroups of order $3^{\ell}$, i.e., $\langle ab\rangle$, $\langle ab^{-1}\rangle$, and
    any of the other two non-characteristic subgroups of $3$, i.e., $\langle a^{3^{\ell-1}}b\rangle$, $\langle a^{-3^{\ell-1}}b\rangle$.
Consider the normalizer $N={\bf N}_{{\sf Aut}(F_{3})}(H_{0})$. For any $g\in N$, $H_{0}=H_{0}^{g}$ stabilizes subgroups $\langle a\rangle^{g}$ and $\langle b\rangle^{g}$, and so $\langle a\rangle^{g}=\langle a\rangle$, $\langle b\rangle^{g}=\langle b\rangle$, which implies that $N\leqslant{\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle)$ is abelian. Note that any subgroup $H\leqslant{\sf Aut}(F_{3})$ with $H\cong{\sf Z}_{2}^{2}$ is a Hall $3^{\prime}$-subgroup, then conjugate to $H_{0}$. Then the normalizer ${\bf N}_{{\sf Aut}(F_{3})}(H)$, which is conjugate to $N$, is also abelian. Thus, ${\sf Aut}(F_{3})$ has no subgroup isomorphic to ${\rm A}_{4}$.
Finally, with Case 3 excluded, the proof is complete. $\Box$
Lemma 5.7.
Let $K$ be a $\{2,3,7\}$-group satisfying Hypotheses 2.2, with neither $K_{3}$ nor $K_{7}$ normal in $K$. Then $K=K_{2}{:}(K_{7}{:}K_{3})={\sf Z}_{2}^{3}{:}(K_{7}{:}K_{3})$.
Proof.
By Lemma 5.1, $K_{\{3,7\}}=K_{7}{:}K_{3}$. Thus $K_{\{2,7\}}\neq K_{7}{:}K_{2}$, and by Lemma 5.6,
$K_{\{2,7\}}=K_{2}{:}K_{7}={\sf Z}_{2}^{3}{:}K_{7},$
Let $F={\rm Fit}(K)=F_{2}\times F_{3}\times F_{7}$. As $K_{2}$ is normalized by $K_{7}$, and meanwhile it normalizes $F_{7}$, so $K_{2}$ centralizes $F_{7}$. Moreover, $K_{2}$ also centralizes $F_{3}$. If not, consider the subgroup $F_{3}{:}(K_{2}{:}K_{7})\leqslant K$. Then ${\sf Z}_{2}^{3}$ as a minimal normal subgroup of $K_{2}{:}K_{7}$ is embedded into ${\sf Aut}(F_{3})$, which is impossible due to Lemma 4.3 and the fact ${\sf Z}_{2}^{3}\nleqslant{\rm GL}(2,3)\cong{\rm Q}_{8}{:}{\rm S}_{3}$. Hence $K_{2}$ centralizes $F$, so $K_{2}\leqslant C_{K}(F)\leqslant F$, and then $K_{2}=F_{2}\lhd K$, and $K=K_{2}{:}K_{\{3,7\}}=K_{2}{:}(K_{7}{:}K_{3})$. $\Box$
We are now ready to prove the first main theorem of this paper.
Proof of Theorem 1.1:
Let $G$ be a finite solvable group of which each Sylow subgroup has a cyclic or dihedral subgroup of prime index. Let $H$ be the largest normal Hall subgroup of odd order, with $\pi=\pi(H)$. Let $K$ be a Hall $\pi^{\prime}$-subgroup of $G$, where $\pi^{\prime}=\pi(G)\setminus\pi(H)$. Then we have factorization
G=H:K.G=H{:}K.
Since $H$ is of odd order, by Lemma 5.4, H=A:BH=A{:}B such that $A$ is abelian, $B$ is nilpotent, and $A\cap{\bf Z}(H)=1$.
We then consider the subgroup $K$. Let $p$ be the largest prime divisor of $|G|$. If $p>7$, then by Lemmas 5.1 and 5.6, $G_{p}$ is normal in both $G_{2^{\prime}}$ and $G_{\{2,p\}}$, and so $G_{p}$ is normal in $G$, and $p\in\pi$. Then, by inductively considering $G/G_{p}$ we deduce that any prime divisor $r$ of $|G|$ with $r>7$ lies in $\pi$. That is, $\pi^{\prime}\subseteq\{2,3,5,7\}$. Further, by Lemmas 5.1, 5.2 and 5.6, $G_{5}$ is normalized by $G_{3}$, $G_{7}$ and $G_{2}$, respectively. Hence $5\notin\pi^{\prime}$. Thus, either
  • (i)
    $\pi^{\prime}=\emptyset$, $G=H$ is of odd order; or
  • (ii)
    $\pi^{\prime}=\{2\}$, $\{2,3\}$, $\{2,7\}$ or $\{2,3,7\}$.
At last, respectively for $\pi^{\prime}=\{2\}$, $\pi^{\prime}=\{2,3\}$ or $\{2,7\}$, and $\pi^{\prime}=\{2,3,7\}$, the group $K=G_{\pi^{\prime}}$ is described in Proposition 4.1, Lemma 5.6 and Lemma 5.7. $\Box$
6. Proof of Theorem 1.2
In this section we focus on the automorphism groups of regular maps.
The following lemma shows that the class of groups with regular triples is closed under taking quotients.
Lemma 6.1.
Let $G$ be a finite group which has a regular triple, and let $\overline{G}$ be a non-identity factor group of $G$. Then either $\overline{G}$ has a regular triple, or $\overline{G}$ is ${\sf Z}_{2}$ or a dihedral group. In particular, $\overline{G}$ is of even order.
Proof.
Let $(x,y,z)$ be a regular triple for $G$, and let $\overline{G}=G/N$ for a proper normal subgroup $N\lhd G$. If $\overline{x},\overline{y},\overline{z}$ are all involutions of $\overline{G}$ with $\overline{x}\neq\overline{z}$, then they form a regular triple for $\overline{G}$, since $\overline{G}=\langle\overline{x},\overline{y},\overline{z}\rangle$ and $\overline{x}\,\overline{z}=\overline{z}\,\overline{x}$; otherwise, either some of $\overline{x},\overline{y},\overline{z}$ equals the identity-element in $\overline{G}$, or $\overline{x}=\overline{z}$, while in both cases, $\overline{G}$ is generated by one or two involutions, namely, $G$ is ${\sf Z}_{2}$ or dihedral, respectively. $\Box$
Lemma 6.2.
Let $G={\rm Q}_{8}{:}{\rm S}_{3}\cong{\rm GL}(2,3)$. Then neither $G$ nor ${\sf Z}_{4}\circ G$ has a regular triple.
Proof.
Suppose that $G$ has a regular triple $(x,y,z)$. Since $xz=zx$, we have $\langle x,z\rangle\cong{\sf Z}_{2}^{2}$, and so $\{x,z,xz\}\cap{\rm Q}_{8}\not$$\emptyset$. That is, one of $x,z,xz$ has to be the central involution 1-1 of $G$. It follows that $\overline{G}=G/\langle-1\rangle=\langle\overline{x},\overline{y}\rangle$ or $\langle\overline{y},\overline{z}\rangle$ is a dihedral group, which is impossible as $\overline{G}={\sf Z}_{2}^{2}{:}{\rm S}_{3}$.
Suppose that $K={\sf Z}_{4}\circ({\rm Q}_{8}{:}{\rm S}_{3})$ has a regular triple $(x,y,z)$. Write $K$ as
$K=\langle u\rangle\circ({\rm Q}_{8}{:}(\langle\sigma\rangle{:}\langle\tau\rangle)),$
where $|u|=4$, ${\rm Q}_{8}=\pm\{1,i,j,k\}$ with $i^{2}=j^{2}=k^{2}=ijk=-1$, and $\sigma,\tau\in{\sf Aut}({\rm Q}_{8})$ such that
$\sigma:i\rightarrow j\rightarrow k\,(\rightarrow i),\mbox{\,\ and\ \,}\tau:i\leftrightarrow-j,j\leftrightarrow-i,k\leftrightarrow-k.$
Then the subgroup $\langle\sigma\rangle{:}\langle\tau\rangle\cong{\rm S}_{3}$ contains three involutions, i.e., $\tau_{k}=\tau$,
$\tau_{i}:j\leftrightarrow-k,k\leftrightarrow-j,i\leftrightarrow-i,\mbox{\ and\ }\tau_{j}:i\leftrightarrow-k,k\leftrightarrow-i,j\leftrightarrow-j.$
By calculation, we conclude that $K$ contains the following involutions:
$\{-1\}\cup\{\pm u\lambda\}\cup\{\pm\tau_{\lambda},\pm\lambda\tau_{\lambda}\},\mbox{\ where\ }\lambda\in\{i,j,k\}.$
We shall confirm the following claim:
Claim. The only involutions commutes with $u\lambda$ (or $u\lambda$) are $\pm u\lambda$ and 1-1.
In fact, for any $\lambda^{\prime}\in\{i,j,k\}$ with $\lambda\neq\lambda^{\prime}$, we have:
$\lambda^{\lambda^{\prime}}=\lambda^{\tau_{\lambda}}=\lambda^{\lambda\tau_{\lambda}}=-\lambda$, and $\lambda^{\tau_{\lambda^{\prime}}}=-\lambda^{\prime\prime},\ \lambda^{\lambda^{\prime}\tau_{\lambda^{\prime}}}=-\lambda^{\tau_{\lambda^{\prime}}}=\lambda^{\prime\prime}$,
where $\lambda^{\prime\prime}\in\{i,j,k\}\setminus\{\lambda,\lambda^{\prime}\}.$ As $u$ and 1-1 lie in the center of $K$, we then have:
$(\pm u\lambda)^{\pm u\lambda^{\prime}}=\pm u(\lambda^{\pm u\lambda^{\prime}})=\pm u(\lambda^{\lambda^{\prime}})=\mp u\lambda,$
$(\pm u\lambda)^{\pm\tau_{\lambda}}=\pm u(\lambda^{\pm\tau_{\lambda}})=\mp u\lambda,\ (\pm u\lambda)^{\pm\lambda\tau_{\lambda}}=\pm u(\lambda^{\pm\lambda\tau_{\lambda}})=\mp u\lambda,$
and
$(\pm u\lambda)^{\pm\tau_{\lambda^{\prime}}}=\pm u(\lambda^{\pm\tau_{\lambda^{\prime}}})=\mp u\lambda^{\prime\prime},\ (\pm u\lambda)^{\pm\lambda^{\prime}\tau_{\lambda^{\prime}}}=\pm u(\lambda^{\pm\lambda^{\prime}\tau_{\lambda^{\prime}}})=\pm u\lambda^{\prime\prime}.$
We have therefore confirmed the claim.
Let $H=\langle u,i,j,k\rangle\leqslant K$. Then $H\cong{\sf Z}_{4}\circ{\rm Q}_{8}$ is normal in $K$, and $K/H\cong{\rm S}_{3}$. Further, the involutions contained in $H$ are: $\{-1\}\cup\{\pm u\lambda\}$. Recall that $(x,y,z)$ is a regular triple for $G$. Thus, $\langle x,z\rangle\cong{\sf Z}_{2}^{2}$, and so $\langle x,z\rangle\cap H>1$, that is,
$\{x,z,xz\}\cap(\{-1\}\cup\{\pm u\lambda\})\not$$\emptyset$.
By the claim given above, one of $x,z,xz$ has to be 1-1. Thus, $\overline{K}=K/\langle-1\rangle=\langle\overline{x},\overline{y}\rangle$ or $\langle\overline{y},\overline{z}\rangle$ is a dihedral group, which is impossible as $\overline{K}=K/\langle-1\rangle={\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}{\rm S}_{3})$. $\Box$
Now we are ready to prove Theorem 1.2.
Proof of Theorem 1.2:
Let $\mathcal{M}$} be a regular map with square-free Euler characteristic, and let $G={\sf Aut}{\mathcal{M}}$. By Lemma 2.1, each Sylow subgroup of $G$ has a cyclic or dihedral subgroup of prime index, and so G=(A:B):KG=(A{:}B){:}K as described in Theorem 1.1.
Note that $G$ has a regular triple. Then it is of even order, and so $K\neq 1$. Further, if $K=G_{\{2,7\}}={\sf Z}_{2}^{3}{:}G_{7}$ or $K=G_{\{2,3,7\}}={\sf Z}_{2}^{3}{:}(G_{7}{:}G_{3})$, then $K$ has $G_{2}={\sf Z}_{2}^{3}$ as a minimal normal subgroup. By checking the $p$-groups appearing in Lemma 4.3, meanwhile noticing that ${\sf Z}_{2}^{3}$ cannot be embedded into ${\rm GL}(2,p)$ for all primes $p$ (as a faithful representation of ${\sf Z}_{2}^{3}$ on $\mathbb{F}_{p}$ is of degree at least $3$), we deduce that $G_{2}={\sf Z}_{2}^{3}$ centralizes each Sylow subgroup of A:BA{:}B. Hence $G_{2}\lhd G$, and so $G$ has an odd factor group $G/G_{2}$, leading to a contradiction by Lemma 6.1. Therefore, $K=G_{2}$ or $G_{\{2,3\}}$.
Suppose $K=G_{2}$. Since K=G/(A:B)K=G/(A{:}B) is a factor group of $G$, by Lemma 6.1, either $K$ has a regular triple and then it is one of the groups given in Proposition 4.2; or $K$ is ${\sf Z}_{2}$ or a dihedral group. In conclusion, $K$ is ${\sf Z}_{2}$, ${\rm D}_{2^{\ell+1}}$, ${\rm D}_{2^{\ell+1}}\times{\sf Z}_{2}$ or ${\rm D}_{2^{\ell+3}}{:}{\sf Z}_{2}$, where $\ell\geqslant 1$.
Suppose $K=G_{\{2,3\}}$, which is one of the groups given in Lemma 5.6 (ii) - (iii). Since K=G/(A:B)K=G/(A{:}B) is a factor group of $G$, by Lemma 6.1, either $K$ has a regular triple, or $K$ is a dihedral group. Further, note that a factor group of a dihedral group is also dihedral (or ${\sf Z}_{2}$). Thus, $K$ has no odd factor groups, and has no factor groups isomorphic to ${\rm Q}_{8}{:}{\rm S}_{3}$ or ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\rm S}_{3})$ either, because by Lemma 6.2 neither ${\rm Q}_{8}{:}{\rm S}_{3}$ nor ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\rm S}_{3})$ has a regular triple, and meanwhile, neither of them is a dihedral group. Then, by checking the groups appearing in Lemma 5.6 (ii) - (iii), we find that the only candidates for $K=G_{\{2,3\}}$ are the following groups in Table 1: $K=X$ or $X\times{\sf Z}_{2}$, where $X\in\{{\rm S}_{4},\,{\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}},\,({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2},\,({\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}\};$ and the two groups in Table 2: $K={\rm D}_{2.3^{\ell}}\times{\rm S}_{4}$ or $({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\rm D}_{6}$. $\Box$
Declarations
Funding and/or Conflicts of interests/Competing interests
The author(s) declares that there is no any financial or personal relationship with other people or organizations not mentioned that can inappropriately influence the work.