### v2

# Finite groups and arc-transitive maps of square-free Euler characteristic

Abstract.
A characterization is completed for finite groups acting arc-transitively on maps with square-free Euler characteristic, associated with infinite families of regular maps of square-free Euler characteristic presented. This is based on a classification of finite groups of which each Sylow subgroup has a cyclic or dihedral subgroup of prime index.
Keywords: solvable group, arc-transitive map, regular map, Euler characteristic
The project was partially supported by the NNSF of China (11931005).
December 1, 2025
1. Introduction
Edge-transitive maps are categorized into fourteen classes according to local structures and local actions in [9, 16], among which five are arc-transitive. The problem of constructing and classifying important classes of such maps with specific Euler characteristic has attracted considerable attention; refer to [4] - [7] for the Euler characteristic to be a negative prime or product of two primes. We aim towards to a classification of arc-transitive maps with square-free Euler characteristic.
Assume that $G$ is a finite group acting edge-transitively on a map with square-free Euler characteristic, which is said to be of square-free Euler characteristic for convenience. Then $G$ has the following restricted property (Lemma 2.1):
each Sylow subgroup has a cyclic or dihedral subgroup of prime index.
It leads us to studying groups of square-free Euler characteristic, which in fact extends the characterization of almost-Sylow cyclic groups, namely, each Sylow subgroup of odd order is cyclic, and the Sylow $2$-subgroup either is trivial or contains a cyclic subgroup of index 2; refer to Zassenhaus [19], Suzuki [17] and Wong [18].
In previous work [11], we characterized non-solvable groups of square-free Euler characteristic, and in [10], we determined $2$-groups of square-free Euler characteristic. As the third paper of this series, the purpose of this paper is to present a classification (characterization) of such groups in the solvable case.
As usual, for a group $G$ and a subset $\pi$ of prime divisors of $|G|$, denote by $G_{\pi}$ a Hall $\pi$-subgroup of $G$, which is a Sylow subgroup if $\pi$ contains a single prime $p$, and denoted by $G_{p}$. For a positive integer $n$, denote by ${\sf Z}_{n}$ a cyclic group of order $n$, by ${\rm D}_{2n}$ or ${\rm Q}_{2n}$ a dihedral or generalized quaternion group of order $2n$, and by ${\rm A}_{n}$ or ${\rm S}_{n}$ an alternating or symmetric group of degree $n$, respectively. A finite group $G$ is a semidirect product of $A$ by $B$ if $A\lhd G$ and $B<G$ are such that $G/A\cong B$, denoted by G=A:BG=A{:}B. Our first main result is the following theorem.
Theorem 1.1.
Let $G$ be a finite solvable group such that each Sylow subgroup of $G$ has a cyclic or dihedral subgroup of prime index. Then:
(1) The Sylow subgroups of $G$ are characterized in Proposition 4.1 (for $G_{2}$) and Lemma 4.3 (for $G_{p}$, $p$ odd prime).
(2) G=(A:B):KG=(A{:}B){:}K, where $\gcd(|A{:}B|,|K|)=1$, such that
  • (i)
    AA is abelian, $B$ is nilpotent, $A\cap{\bf Z}(A{:}B)=1;$
  • (ii)
    $K$ is one of the following groups:
    • (a)
      $K=G_{2}$ ($K$ might be $1$ in case of $G$ of odd order);
    • (b)
      $K=G_{\{2,3\}}$, explicitly given in Lemma 5.6 (2) - (3);
    • (c)
      $K=G_{\{2,7\}}={\sf Z}_{2}^{3}{:}G_{7}$, or $K=G_{\{2,3,7\}}={\sf Z}_{2}^{3}{:}(G_{7}{:}G_{3})$.
The next theorem shows that arc-transitive automorphism groups of maps with square-free Euler characteristic are more restricted.
Theorem 1.2.
Let $\mathcal{M}$} be a map with square-free Euler characteristic, and let $G\leqslant{\sf Aut}{\mathcal{M}}$ be solvable and arc-transitive on $\mathcal{M}$}. Then G=(A:B):KG=(A{:}B){:}K such that $A$ is abelian with $A\cap{\bf Z}(A{:}B)=1$, $B$ is nilpotent, and either $K=G_{2}$ is ${\sf Z}_{2}$ or one of the groups given in Proposition 4.2, or one of the following holds:
  • (1)
    $\mathcal{M}$} is $G$-regular, $K=G_{\{2,3\}}$ is one of $X$, $X\times{\sf Z}_{2}$, ${\rm D}_{2.3^{\ell}}\times{\rm S}_{4}$ or $({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\rm D}_{6}$, where $X\in\{{\rm S}_{4},\,{\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}},\,({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2},\,({\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}\};$
  • (2)
    $\mathcal{M}$} is $G$-vertex-reversing, $K=G_{\{2,3\}}$ is one of the groups in Table 1 and Table 2, corresponding to cases (1.1), (1.2), (1.5), (1.6), (2.4), (2.5);
  • (3)
    $\mathcal{M}$} is $G$-vertex-rotary, $K=G_{\{2,7\}}={\sf Z}_{2}^{3}{:}{\sf Z}_{7^{\ell}}$, or $K=G_{\{2,3\}}$ is one of ${\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}}$, ${\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})$, ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})$, or one of the groups in Table 1 and Table 2, excluding cases (1.5), (1.6).
The proofs of the above two results are given in Section 4 - 5 and Section 6. Before that, we construct several interesting examples of regular maps with square-free Euler characteristic in Section 3, from which the following result follows.
Theorem 1.3.
There are infinite families of regular maps $\mathcal{M}$} with underlying graphs ${\bf C}_{n}^{(n)}$ or ${\bf C}_{n}\square{\bf C}_{n}$, and with Euler characteristic $\chi({\mathcal{M}})=-n(n-3)$ or n(n2)-n(n-2), respectively. In particular, there are infinitely many of them with $\chi({\mathcal{M}})$ square-free.
2. Preliminaries
Let $G$ be a group. Denote by $\Phi(G)$ and $G^{\prime}$ the Frattini subgroup and the commutator subgroup of $G$. If $G$ is a $p$-group, then $^{\prime}\leqslant\Phi(G), G/G$$G^{\prime}$ is abelian, and $G/\Phi(G)$ is elementary abelian. For a subgroup $H\leqslant G$, denote by ${\bf N}_{G}(H)$ and ${\bf C}_{G}(H)$ the normalizer and the centralizer of $H$ in $G$, respectively.
Let ${\it\Gamma},\Sigma$ be two graphs and $\lambda$ be a positive integer. Denote by ${\it\Gamma}^{(\lambda)}$ the graph with each edge of $\it\Gamma$} replaced by $\lambda$ edges. Denote by ${\it\Gamma}\square\Sigma$ the Cartesian product of $\it\Gamma$} and $\Sigma$, which has vertex set $V({\it\Gamma})\times V(\Sigma)$, where $(x_{1},y_{1})$ is adjacent to $(x_{2},y_{2})$ if and only if $x_{1}=y_{1}$ and $x_{2}$ is adjacent to $y_{2}$, or $x_{1}$ is adjacent to $y_{1}$ and $x_{2}=y_{2}$.
A map ${\mathcal{M}}=(V,E,F)$ is a $2$-cell embedding of a graph into a closed surface, with vertex set $V$, edge set $E$ and face set $F$. The Euler characteristic of $\mathcal{M}$} is given by Euler formula:
$\chi({\mathcal{M}})=|V|-|E|+|F|.$
A flag of $\mathcal{M}$} is an incidence triple $(\alpha,e,f)$ with $\alpha\in V$, $e\in E$ and $f\in F$. Let $G\leqslant{\sf Aut}({\mathcal{M}})$, and let $(\alpha,e,f)$ be a flag. It is well-known that each of the stabilizers $G_{\alpha}$ and $G_{f}$ is cyclic or dihedral, and meanwhile, $G_{e}=1$, ${\sf Z}_{2}$ or ${\rm D}_{4}$.
A map $\mathcal{M}$} is called a $G$-arc-transitive map, if $G$ acts transitively on arcs (directed edges) of $\mathcal{M}$}. Arc-transitive maps are divided into five types in [9, 16], for which recent papers [12, 13] provide a new description: letting $\mathcal{M}$} be a $G$-arc-transitive map, then one of the following holds:
(1) $\mathcal{M}$} is a $G$-regular map (type 1), which is determined a regular triple $(x,y,z)$ for $G$ satisfying:
$\mbox{$G=\langle x,y,z\rangle$, $|x|=|y|=|z|=2$ and $\langle x,z\rangle\cong{\sf Z}_{2}^{2}$}.$
In this case, we have:
$G={\sf Aut}({\mathcal{M}}),\,G_{\alpha}=\langle x,y\rangle,\,G_{e}=\langle x,z\rangle,\,G_{f}=\langle y,z\rangle,$
and the underlying graph of $\mathcal{M}$} can be expressed as a coset graph ${\sf Cos}(G,H,HzH)$, where $H=\langle x,y\rangle$ (refer to [12].)
(2) $\mathcal{M}$} is a $G$-vertex-reversing map (type $2^{\ast}$ or $2^{P}$), which is determined by a reversing triple $(x,y,z)$ for $G$ satisfying: $G=\langle x,y,z\rangle$ and $|x|=|y|=|z|=2$.
(3) $\mathcal{M}$} is a $G$-vertex-rotary map (type $2^{*}{\rm ex}$ or $2^{P}{\rm ex}$), which is determined by a rotary pair $(\alpha,z)$ for $G$ satisfying: $G=\langle\alpha,z\rangle$ and $|z|=2$.
The following lemma established in [11] characterizes Sylow subgroups of ${\sf Aut}({\mathcal{M}})$ in the case where $\mathcal{M}$} is with square-free characteristic. We note that if $\chi({\mathcal{M}})$ is square-free, then of course so is $\gcd(\chi({\mathcal{M}}),|{\sf Aut}{\mathcal{M}}|)$.
Lemma 2.1.
([11, Lem. 2.2]) Let $\mathcal{M}$} be a map such that $\gcd(\chi({\mathcal{M}}),|{\sf Aut}({\mathcal{M}})|)$ is square-free. Then each Sylow $p$-subgroup of ${\sf Aut}({\mathcal{M}})$ has a cyclic or dihedral subgroup of index $p$.
We thus study finite groups satisfying the following hypothesis.
Hypothesis 2.2.
Let $G$ be a finite group such that each Sylow subgroup of $G$ has a cyclic or dihedral subgroup of prime index.
The next lemma plays a key role in the analysis of groups satisfying Hypothesis 2.2.
Lemma 2.3.
([11, Lem. 2.5]) Let a finite group $G$ satisfy Hypothesis 2.2. Then so does each subgroup and each factor group of $G$.
3. Examples of square-free Euler characteristic regular maps
We construct several infinite families of regular maps with square-free Euler characteristic in this section. The constructions are based on the following groups:
$\begin{array}[]{lllll}X&=&(\langle a,s\rangle\times\langle b,t\rangle){:}\langle\sigma\rangle\cong{\rm D}_{2n}\wr{\rm S}_{2}&&\\ &=&\langle a,b\rangle{:}\langle s,\sigma\rangle\cong{\sf Z}_{n}^{2}{:}{\rm D}_{8},&&(\ast)\end{array}$
where $\langle a,s\rangle=\langle a\rangle{:}\langle s\rangle={\rm D}_{2n}$, $\langle b,t\rangle=\langle b\rangle{:}\langle t\rangle={\rm D}_{2n}$, $(a,s)^{\sigma}=(b,t)$ and $|\sigma|=2$.
The first family is given now.
Construction 3.1.
Let $G=\langle a,s\rangle\times\langle b,t\rangle={\rm D}_{2n}\times{\rm D}_{2n}\leqslant X$ as given in ($\ast$), with $n$ odd. Let
x=s,y=abst,z=st.x=s,\ y=abst,\ z=st.
Then $(x,y,z)$ is a triple of involutions with $xz=zx$. Further,
$(yx)^{2}=(abt)^{2}=a^{2}$ and $(yxz)^{2}=(abs)^{2}=b^{2}$.
As $n$ is odd, $\langle a\rangle=\langle a^{2}\rangle$ and $\langle b\rangle=\langle b^{2}\rangle$, so that $G=\langle x,y,z\rangle$. Thus, $(x,y,z)$ is a regular triple for $G$.
Lemma 3.2.
Let ${\mathcal{M}}_{n}={\mathcal{M}}(G,x,y,z)$ be a regular map defined in Construction 3.1. Then the underlying graph of ${\mathcal{M}}_{n}$ is a multiple cycle ${\bf C}_{n}^{(n)}$, and $\chi({\mathcal{M}}_{n})=-n(n-3)$. In particular, $\chi({\mathcal{M}}_{n})$ is square-free for infinitely many odd values of $n$.
Proof.
The vertex stabilizer $H$ is
$H=\langle x,y\rangle=\langle s,abst\rangle=\langle abt\rangle{:}\langle s\rangle=\langle a,s\rangle\times\langle bt\rangle={\rm D}_{4n}.$
Meanwhile, the face stabilizer is
$\langle y,z\rangle=\langle abst,st\rangle=\langle ab\rangle{:}\langle st\rangle={\rm D}_{2n}.$
Thus the map ${\mathcal{M}}_{n}$ has valency $2n$ and face length $n$. Further, $H\cap H^{z}=\langle a,s\rangle\cong{\rm D}_{2n}$ is of index $2$ in $H$, so the underlying graph ${\it\Gamma}={\sf Cos}(G,H,HzH)\cong{\bf C}_{n}^{(n)}$.
The Euler characteristic
$\chi({\mathcal{M}}_{n})=|G|\left({1\over|\langle x,y\rangle|}-{1\over 4}+{1\over|\langle y,z\rangle|}\right)=4n^{2}\left({1\over 4n}-{1\over 4}+{1\over 2n}\right)=-n(n-3).$
Let $n=4x+1$. Then $\chi({\mathcal{M}}_{n})=-2(4x+1)(2x-1)$. By a result given by G. Ricci in 1933 (which is improved by $\rm P.Erd\ddot{o}s$} in [8]), there are infinitely many positive integers $x$ for which f(x)=(4x+1)(2x1)f(x)=(4x+1)(2x-1) is square-free, so there are infinitely many odd values of $n$ such that $\chi({\mathcal{M}}_{n})$ is square-free. $\Box$
Here is some data such that n(n3)-n(n-3) is square-free, with $n$ odd.
$\begin{array}[]{c|ccccccc}\hline\cr n&5&13&17&29&37&41&\cdots\\ -n(n-3)&-2.5&-2.5.13&-2.7.17&-2.13.29&-2.17.37&-2.19.41$$\cdots\\ \hline\cr\end{array}$
Next we construct the second family.
Construction 3.3.
Let $G=\langle a,b\rangle{:}\langle st,\sigma\rangle={\sf Z}_{n}^{2}{:}{\sf Z}_{2}^{2}\leqslant X$ as given in ($\ast$). Let
$x=\sigma,\ y=ast,\ z=abst.$
Then $(x,y,z)$ is a triple of involutions with $xz=zx$. Further,
$yz=b^{-1}$ and $(yz)^{x}=a^{-1}$,
so that $G=\langle x,y,z\rangle$. Thus, $(x,y,z)$ is a regular triple for $G$.
Lemma 3.4.
Let ${\mathcal{M}}_{n}={\mathcal{M}}(G,x,y,z)$ be a regular map defined in Construction 3.3. Then the underlying graph of ${\mathcal{M}}_{n}$ is ${\bf C}_{n}^{(n)}$, and $\chi({\mathcal{M}}_{n})=-n(n-3)$. In particular, $\chi({\mathcal{M}}_{n})$ is square-free for infinitely many values of $n$.
Proof.
The vertex stabilizer $H$ is
$H=\langle x,y\rangle=\langle ast\sigma\rangle{:}\langle\sigma\rangle={\rm D}_{4n}.$
Meanwhile, the face stabilizer is
$\langle y,z\rangle=\langle abst,ast\rangle=\langle b\rangle{:}\langle ast\rangle={\rm D}_{2n}.$
Thus the map ${\mathcal{M}}_{n}$ has valency $2n$ and face length $n$. Further, $H\cap H^{z}=\langle a^{-1}b\rangle{:}\langle\sigma\rangle={\rm D}_{2n}$ is of index $2$ in $H$, so the underlying graph ${\it\Gamma}={\sf Cos}(G,H,HzH)\cong{\bf C}_{n}^{(n)}$.
The Euler characteristic
$\chi({\mathcal{M}})=|G|\left({1\over|\langle x,y\rangle|}-{1\over 4}+{1\over|\langle y,z\rangle|}\right)=4n^{2}\left({1\over 4n}-{1\over 4}+{1\over 2n}\right)=-n(n-3).$
By Lemma 3.2, there are infinitely many odd values of $n$ such that $\chi({\mathcal{M}}_{n})$ is square-free. Further, let $n=4x+2$. Then according to [8], there are infinitely many positive integers $x$ for which f(x)=(2x+1)(4x1)f(x)=(2x+1)(4x-1) is square-free, so there are infinitely many even values of $n$ such that $\chi({\mathcal{M}}_{n})=-2(2x+1)(4x-1)$ is square-free. $\Box$
Here is some data such that n(n3)-n(n-3) is square-free, with $n$ even.
$\begin{array}[]{c|ccccccc}\hline\cr n&2&10&14&22&26&34&\cdots\\ -n(n-3)&2&-2.5.7&-2.7.11&-2.11.19&-2.13.23&-2.17.31$$\cdots\\ \hline\cr\end{array}$
The third family is given below.
Construction 3.5.
Let $G=\langle a,b\rangle{:}\langle s,\sigma\rangle={\sf Z}_{n}^{2}{:}{\rm D}_{8}=X$ as given in ($\ast$), with $n$ odd. Let
$x=\sigma,\ y=s,\ z=abst.$
Then $(x,y,z)$ is a triple of involutions with $xz=zx$. Further,
$(zy)^{2}=(abt)^{2}=a^{2}$ and $((zy)^{2})^{x}=(a^{2})^{x}=b^{2}$.
As $n$ is odd, $\langle a\rangle=\langle a^{2}\rangle$ and $\langle b\rangle=\langle b^{2}\rangle$, so that $G=\langle x,y,z\rangle$. Thus, $(x,y,z)$ is a regular triple for $G$.
Lemma 3.6.
Let ${\mathcal{M}}_{n}={\mathcal{M}}(G,x,y,z)$ be a regular map defined in Construction 3.5. Then the underlying graph of ${\mathcal{M}}_{n}$ is ${\bf C}_{n}\square{\bf C}_{n}$, and $\chi({\mathcal{M}}_{n})=-n(n-2)$. In particular, $\chi({\mathcal{M}}_{n})$ is square-free for infinitely many prime values of $n$.
Proof.
The vertex stabilizer $H$ is
$H=\langle x,y\rangle=\langle\sigma,s\rangle={\rm D}_{8}.$
Meanwhile, the face stabilizer
$\langle y,z\rangle=\langle s,abst\rangle=\langle abt\rangle{:}\langle s\rangle={\rm D}_{4n}.$
Thus the map ${\mathcal{M}}_{n}$ has valency $4$ and face length $2n$. Let ${\it\Gamma}={\sf Cos}(G,H,HzH)$ be the underlying graph of ${\mathcal{M}}_{n}$. Then $G=\langle a,b\rangle{:}H$, and so $\langle a,b\rangle$ is regular on the vertex set of $\it\Gamma$}. Hence $\it\Gamma$} is a (connected and undirected) Cayley graph ${\sf Cay}(K,S)$ of valency 4, where $K=\langle a,b\rangle={\sf Z}_{n}^{2}$. It follows that $S=\{x,x^{-1},y,y^{-1}\}$ with $\langle x,y\rangle=\langle a,b\rangle$, and so ${\it\Gamma}={\bf C}_{n}\square{\bf C}_{n}$.
The Euler characteristic
$\chi({\mathcal{M}}_{n})=|G|\left({1\over|\langle x,y\rangle|}-{1\over 4}+{1\over|\langle y,z\rangle|}\right)=8n^{2}\left({1\over 8}-{1\over 4}+{1\over 4n}\right)={-n(n-2).}$
According to [8], there are infinitely many odd values of $n$ such that $\chi({\mathcal{M}}_{n})$ is square-free. Further, by a theorem of Chen [3], there indeed exist infinitely many primes $p$ such that p2p-2 is square-free, then so is p(p2)-p(p-2). $\Box$
Here is some data such that p(p2)-p(p-2) is square-free, with $p$ prime.
$\begin{array}[]{c|cccccccc}\hline\cr p&3&5&7&13&17&19&23&\cdots\\ -p(p-2)&-3&-3.5&-5.7&-11.13&-3.5.17&-17.19&-3.7.23$$\cdots\\ \hline\cr\end{array}$
4. $p$-Groups
In this section, we study $p$-groups satisfying Hypothesis 2.2, i.e., $p$-groups with a cyclic or dihedral subgroup of index $p$. Note that the case for $p=2$ is dealt with in the previous work [10], which is introduced below.
Let $A,B$ be finite groups whose centers ${\bf Z}(A),{\bf Z}(B)$ have isomorphic subgroups $C_{1},C_{2}$, respectively. Let $\varphi$ be an isomorphism from $C_{1}$ to $C_{2}$, and let $C=\{(c,c^{\varphi})\mid c\in C_{1}\}$. Then $C$ is a normal subgroup of $A\times B$. The factor group $(A\times B)/C$ is called a central product of $A$ and $B$, denoted by $A\circ_{C}B$, and sometimes simply by $A\circ B$ if $C$ is equal to ${\bf Z}(A)$ or ${\bf Z}(B)$.
Let ${\rm Q}_{2^{\ell+1}}=\langle u,v\rangle$ be a generalized quaternion group of order $2^{\ell+1}$, where
$|u|=2^{\ell}\geqslant 2^{2}$, $u^{2^{\ell-1}}=v^{2}$, and $u^{v}=u^{-1}$.
Then ${\rm Q}_{2^{\ell+1}}$ has $u^{2^{\ell-1}}=v^{2}$ as the only involution. Let
$G={\rm Q}_{2^{\ell+1}}\circ{\sf Z}_{4}=\langle u,v\rangle\circ\langle b\rangle=(\langle u,v\rangle\times\langle b\rangle)/\langle b^{2}v^{2}\rangle.$
Then $G$ contains normal subgroups $\langle u,v\rangle={\rm Q}_{2^{\ell+1}}$ and $\langle b\rangle={\sf Z}_{4}$, and the only involutions contained in $\langle u,v\rangle$ and $\langle b\rangle$ are $u^{2^{\ell-1}}=v^{2}$ and $b^{2}$, respectively, which are actually identical in $G$ as $b^{2}v^{2}=1$.
Proposition 4.1.
([10], Thm. 1.1) Let $G$ be a $2$-group which has a cyclic or dihedral subgroup of index $2$. Then $G$ is one of the following groups. Further, either ${\sf Aut}(G)$ is a $2$-group, or $(G,{\sf Aut}(G))$ is one of $({\sf Z}_{2}^{2},{\rm S}_{3}),\,({\sf Z}_{2}^{3},{\rm GL}_{3}(2)),\,({\rm Q}_{8},{\rm S}_{4}),\,({\rm Q}_{8}\circ{\sf Z}_{4},{\rm S}_{4}\times{\sf Z}_{2})$.
  • (1)
    $G$ is one of the following groups, with a cyclic subgroup of index $2$, $\ell\geqslant 1$:
    • (1.1)
      $G={\sf Z}_{2^{\ell+1}}$ or ${\sf Z}_{2^{\ell}}\times{\sf Z}_{2};$
    • (1.2)
      $G={\rm D}_{2^{\ell+2}}$ or ${\rm Q}_{2^{\ell+2}};$
    • (1.3)
      $G={\sf Z}_{2^{\ell+2}}{:}{\sf Z}_{2}=\langle a\rangle{:}\langle b\rangle$ with $a^{b}=a^{2^{\ell+1}\pm 1};$
  • (2)
    $G$ is one of the following groups, with a dihedral subgroup of index $2$, $\ell\geqslant 1$:
    • (2.1)
      $G={\rm D}_{2^{\ell+1}}\times{\sf Z}_{2};$
    • (2.2)
      $G={\rm D}_{2^{\ell+3}}{:}{\sf Z}_{2}=\langle a,b\rangle{:}\langle c\rangle$, where $\langle a\rangle{:}\langle b\rangle={\rm D}_{2^{\ell+3}}$, $(a,b)^{c}=(a^{2^{\ell+1}+1},b);$
    • (2.3)
      $G={\rm Q}_{2^{\ell+2}}\circ{\sf Z}_{4}$.
Proposition 4.2.
([10], Thm. 1.1) Let $G$ be a $2$-group which has a cyclic or dihedral subgroup of index $2$. Then:
  • (1)
    If $G$ has a reversing triple, then $G={\rm D}_{2^{\ell+1}}$, ${\rm D}_{2^{\ell+1}}\times{\sf Z}_{2}$, ${\rm D}_{2^{\ell+3}}{:}{\sf Z}_{2}$ or ${\rm Q}_{2^{\ell+2}}\circ{\sf Z}_{4}$, where $\ell\geqslant 1;$ further, if $G$ has a regular triple, then $G\neq{\rm Q}_{2^{\ell+2}}\circ{\sf Z}_{4};$
  • (2)
    If $G$ has a rotary pair, then $G={\sf Z}_{2^{\ell+1}}$, ${\sf Z}_{2^{\ell}}\times{\sf Z}_{2}$, ${\rm D}_{2^{\ell+2}}$ or ${\sf Z}_{2^{\ell+2}}{:}{\sf Z}_{2}$, $\ell\geqslant 1$.
For the odd primes $p$, we have the following characterization.
Lemma 4.3.
Let $p>2$ be an odd prime, and let $G$ be a $p$-group which has a cyclic subgroup of index $p$. Then one of the following holds. In particular, each prime divisor of $|{\sf Aut}(G)|$ is less than or equal to $p$.
  • (1)
    $G={\sf Z}_{p^{\ell}}$ is cyclic, and ${\sf Aut}(G)={\sf Z}_{(p-1)p^{\ell-1}}$.
  • (2)
    $G={\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$ is abelian, and
    1. (2.1)
      if $\ell>1$, then ${\sf Aut}(G)$ is solvable, and letting $G=\langle a,b\rangle$, then any Hall $p^{\prime}$-subgroup of ${\sf Aut}(G)$ is conjugate to $({\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle))_{p^{\prime}}={\sf Z}_{p-1}\times{\sf Z}_{p-1};$
    2. (2.2)
      if $\ell=1$, then ${\sf Aut}(G)={\rm GL}(2,p)$, and further, any subgroup $M\leqslant{\rm GL}(2,p)$ of order coprime to $2p$ is an abelian group.
  • (3)
    $G=\langle a\rangle{:}\langle b\rangle={\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}$ is non-abelian with $a^{b}=a^{p^{\ell-1}+1}$, $\ell\geqslant 2$, and ${\sf Aut}(G)$ is solvable such that any Hall $p^{\prime}$-subgroups of ${\sf Aut}(G)$ is conjugate to $\langle\rho\rangle={\sf Z}_{p-1}$, where $a^{\rho}=a^{\xi}\not=a$ with $\xi\,\,\big|\,\,p-1$, and $b^{\rho}=b$.
Proof.
According to [14, 5.3.4, pp.136], $G$ is one of the groups stated in this lemma. We only need to verify the statements regarding the automorphism groups of $G$.
(i) The automorphism groups of cyclic groups are well-known; refer to [15, Thm. 7.3].
(ii) Suppose $G={\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$ with $\ell>1$. Then by [1, Thm. 3.2], ${\sf Aut}(G)$ is of order
$\begin{array}[]{ccl}|{\sf Aut}(G)|&=&|{\sf Aut}({\sf Z}_{p^{\ell}})|\cdot|{\sf Aut}({\sf Z}_{p})|\cdot|{\sf Hom}({\sf Z}_{p^{\ell}},{\sf Z}_{p})|\cdot|{\sf Hom}({\sf Z}_{p},{\sf Z}_{p^{\ell}})|\\ &=&(p-1)p^{\ell-1}\cdot(p-1)\cdot p\cdot p=p^{\ell+1}(p-1)^{2}.\end{array}$
More explicitly, let $G=\langle a\rangle\times\langle b\rangle$ with $|a|=p^{\ell}$, $|b|=p$. Then $\Phi(G)=\langle a^{p}\rangle$. Let $\overline{G}=G/\Phi(G)$, and consider the homomorphism
$\sigma:{\sf Aut}(G)\rightarrow{\sf Aut}(\overline{G})={\rm GL}(2,p).$
On the one hand, the kernel of $\sigma$ consists of automorphisms $g\in{\sf Aut}(G)$ which stabilizes each of the $\Phi(G)$-cosets, and in particular, stabilizes $\Phi(G)\langle a\rangle$ and $\Phi(G)\langle b\rangle$. Thus, by direct calculation we have:
$a^{g}\in\{a,a^{p+1},a^{2p+1},\cdots,a^{(p^{\ell-1}-1)p+1}\}\mbox{\ and\ }b^{g}\in\{b,a^{p^{\ell-1}}b,a^{2p^{\ell-1}}b,\cdots,a^{(p-1)p^{\ell-1}}b\}.$
Consequently, ${\rm Ker}(\sigma)=\langle x\rangle\times\langle y\rangle\cong{\sf Z}_{p^{\ell-1}}\times{\sf Z}_{p},$ where
$x:a\rightarrow a^{p+1},b\rightarrow b\ \mbox{\,and\,}\ y:a\rightarrow a,b\rightarrow a^{p^{\ell-1}}b.$
On the other hand, the image of $\sigma$ falls into the (solvable) parabolic subgroup $P$ of ${\sf Aut}(\overline{G})={\rm GL}(2,p)$ which stabilizes $\langle\overline{b}\rangle$, because
$\langle\overline{b}\rangle=\{\overline{1},\overline{b},\cdots,\overline{b}^{p-1}\}=\{\Phi(G),b\Phi(G),\cdots,b^{p-1}\Phi(G)\}$
exactly consists of those elements of $G$ of order less than $p^{\ell}$. Moreover, it follows from ${\sf Aut}(G)/{\rm Ker}(\sigma)\cong{\rm Im}(\sigma)$ that $|{\rm Im}(\sigma)|=p(p-1)^{2}=|P|$, and so ${\rm Im}(\sigma)=P$.
Now both ${\rm Ker}(\sigma)$ and ${\rm Im}(\sigma)$ are solvable, so is ${\sf Aut}(G)$. Note that the following subgroup
$X=({\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle))_{p^{\prime}}={\sf Z}_{p-1}\times{\sf Z}_{p-1}$
is a Hall $p^{\prime}$-subgroup of $G$. Then any other Hall $p^{\prime}$-subgroup is conjugate to $X$.
(iii) Suppose $G={\sf Z}_{p}\times{\sf Z}_{p}$. Then ${\sf Aut}(G)={\rm GL}(2,p)$ is of order $p(p+1)(p-1)^{2}$, with prime divisors less than or equal to $p$. Further, let $M\leqslant{\rm GL}(2,p)$ be a subgroup of order coprime to $2p$, and let $N=M/{\bf Z}(M)$. Then
$N=M/{\bf Z}(M)\leqslant M/({\bf Z}({\rm GL}(2,p))\cap M)\lesssim{\rm PGL}(2,p).$
Further, since ${\rm PSL}(2,p)N\leqslant{\rm PGL}(2,p)$, we have
$|N/({\rm PSL}(2,p)\cap N)|\leqslant|{\rm PGL}(2,p)|/|{\rm PSL}(2,p)|=2.$
It then follows by $2\nmid|N|$ that $N$ is indeed embedded into ${\rm PSL}(2,p)$. Note that the subgroups of ${\rm PSL}(2,p)$ are all explicitly known, by check which a subgroup of order coprime to $2p$ must be a cyclic group. Then $M$ is an abelian group, since $N=M/{\bf Z}(M)$ is cyclic.
(iv) Suppose that $G=\langle a\rangle{:}\langle b\rangle={\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}$ with $a^{b}=a^{p^{\ell-1}+1}$ is non-abelian, $\ell\geqslant 2$. The automorphism group ${\sf Aut}(G)$ is described in [2, Thm. 3.1 & 3.2], by which we have: $|{\sf Aut}(G)|=(p-1)p^{\ell+1}$, ${\sf Aut}(G)$ is solvable with the Sylow $p$-subgroup normal, and $X=\langle\rho\rangle={\sf Z}_{p-1}$ is a Hall $p^{\prime}$-subgroup of $G$, where $\rho$ stabilizes $\langle a\rangle$ and centralizes $\langle b\rangle$, that is, $a^{\rho}=a^{\xi}\not=a$ with $\xi\,\,\big|\,\,p-1$, and $b^{\rho}=b$. $\Box$
5. Proof of Theorem 1.1
In this section, we study the way how the $p$-groups described in Section 4 compose a larger solvable group, and give a proof of Theorem 1.1
Let $G$ be a finite group, with $\{p_{1},\ldots,p_{s}\}$ the prime divisors of $|G|$. For $p=p_{i}$, denote by ${\bf O}_{p}(G)$ the maximal normal $p$-subgroup of $G$, or equivalently, the intersection of all Sylow $p$-subgroups of $G$. Let $F={\rm Fit}(G)$ be the Fitting subgroup of $G$, which is the largest nilpotent normal subgroup. Then
$F=F_{p_{1}}\times\cdots\times F_{p_{s}},\mbox{\ with\ each\ }F_{p_{i}}={\bf O}_{p_{i}}(G).$
Note that both ${\bf O}_{p}(G)$ and $F$ are characteristic in $G$. Further, if $G$ is solvable, then ${\bf C}_{G}(F)\leqslant F$, and so
$G/F={\bf N}_{G}(F)/\left(F{\bf C}_{G}(F)\right)\lesssim{\sf Out}(F).$
We first consider solvable groups of odd order.
Lemma 5.1.
Let $G$ be a solvable group of odd order satisfying Hypothesis 2.2. Let $p$ be the largest prime divisor of $|G|$. Then a Sylow $p$-subgroup $G_{p}$ is normal in $G$, and so $G=G_{p}{:}G_{p^{\prime}}$.
Proof.
Let $\overline{G}=G/{\bf O}_{p}(G)$, and let $F={\rm Fit}(\overline{G})$. Notice that ${\bf O}_{p}(\overline{G})=1$. We have $F=F_{p_{1}}\times\cdots\times F_{p_{s}}$ with all $p_{i}<p$. Now,
$\overline{G}/F\lesssim{\sf Out}(F)={\sf Out}(F_{p_{1}})\times\cdots\times{\sf Out}(F_{p_{s}}),$
and so $p\nmid|\overline{G}/F|$ by Lemma 4.3. Thus, $p\nmid|\overline{G}|$, so that, $G_{p}={\bf O}_{p}(G)\lhd G$. $\Box$
Lemma 5.2.
Let $G$ be a $\{p,q\}$-group satisfying Hypotheses 2.2, where $p>q$ are both odd primes, and $q\nmid p^{2}-1$. Then $G=G_{p}\times G_{q}$.
Proof.
By Lemma 5.1, $G=G_{p}{:}G_{q}$. Since $G$ is solvable, we obtain that $G_{q}/{\bf C}_{G_{q}}(G_{p})\cong G/(G_{p}{\bf C}_{G_{q}}(G_{p}))=G/(G_{p}{\bf C}_{G}(G_{p}))\leq{\sf Out}(G_{p})$. By Lemma 4.3, any prime divisor of $|{\sf Out}(G_{p})|$ divides $p(p^{2}-1)$. However, $q\nmid p(p^{2}-1)$ by our assumption, and we conclude that $G_{q}/{\bf C}_{G_{q}}(G_{p})=1$, so that $G_{q}={\bf C}_{G_{q}}(G_{p})$ and $G=G_{p}\times G_{q}$. $\Box$
Before giving a general characterization of solvable groups of odd order, we need the following technical lemma.
Lemma 5.3.
Let $H={\sf Z}_{p^{\ell}}$ be a cyclic group and $x$ be an automorphism of $H$ with order coprime to $p$. Let $X=H{:}\langle x\rangle$. Then, (i) $H\leqslant X^{\prime};$ (ii) $H\cap{\bf Z}(X)=1$.
Proof.
By Lemma 4.3 (1), the order of $x$ divides p1p-1. Let $H=\langle a\rangle$ and suppose
$a^{x}=a^{m}$, where $1<m\leqslant p^{\ell}-1$,  $p\nmid m$  and  $p^{\ell}\mid m^{p-1}-1$.
If $p\mid m-1$, then by LTE (Lift the Exponent Lemma) we have $p^{\ell}\mid m-1$, a contradiction. Hence $p\nmid m-1$. Thus, (i) the commutator $a,x]=a^{-1}a^{x}=a^{m-1}$ can generate $H$, so that, $H\leqslant X^{\prime}$; (ii) for any $1\leqslant k<p^{\ell}$, $(a^{k})^{x}=a^{km}\neq a^{k}$, so that, $H\cap{\bf Z}(X)=1$. $\Box$
Lemma 5.4.
Let $G$ be a solvable group of odd order satisfying Hypothesis 2.2. Then G=C:DG=C{:}D, where $C$ is abelian, $D$ is nilpotent, and $C\cap{\bf Z}(G)=1$.
Proof.
Let $n$ be the number of prime divisors of $|G|$. The proof proceeds by induction on $n$. For $n=1$, the lemma holds with $C=1$, $D=G$. Suppose $n\geqslant 2$, and assume inductively that the lemma holds for groups of order with less prime divisors.
Let $p$ be the largest prime divisor of $|G|$. By Lemma 5.1 and the induction assumption, we have
$G=G_{p}{:}G_{p^{\prime}}=G_{p}{:}(A{:}B),$
with $A$ abelian, $B$ nilpotent, and $A\cap{\bf Z}(A{:}B)=1$, so that, $B$ does not centralize any Sylow subgroup of $A$. Further, if $G=G_{p}\times G_{p^{\prime}}$, then
$G=G_{p}{\times}(A{:}B)=(G_{p}{\times}A){:}B=A{:}(G_{p}\times B),$
which gives an expected factorization of $G$. We thus assume that $G\not=G_{p}\times G_{p^{\prime}}$, namely, $G_{p^{\prime}}$ acts by conjugation on $G_{p}$ non-trivially. We shall confirm
Claim: $G_{p}$ is centralized by each Sylow $r$-subgroup $A_{r}$ of $A$, so that, $G_{p}$ is centralized by $A$.
By Lemma 4.3, we have two cases to treat: $A_{r}={\sf Z}_{r^{k}}$ or ${\sf Z}_{r^{k}}\times{\sf Z}_{r}$.
Case 1. Let $A_{r}={\sf Z}_{r^{k}}$. If $A_{r},B_{r^{\prime}}]=1$, then by Lemma 4.3, $(A{:}B)_{r}=A_{r}{:}B_{r}={\sf Z}_{r^{k}}{:}{\sf Z}_{r}$ is non-abelian, and
$1\not=A_{r}\cap{\bf Z}(A_{r}{:}B_{r})=A_{r}\cap{\bf Z}(A{:}B)\leqslant A\cap{\bf Z}(A{:}B),$
a contradiction. Thus there exists a Sylow $q$-subgroup $B_{q}$ with $q\not=r$ which does not centralize $A_{r}$. Take $x\in B_{q}$ to be an element acting non-trivially on $A_{r}$ and let $X=A_{r}{:}\langle x\rangle$. Then $X$ acts on $G_{p}$ and induces an abelian automorphism group of $G_{p}$ by Lemma 4.3. It then follows by Lemma 5.3 (i) that $A_{r}\leqslant X^{\prime}\leqslant{\bf C}_{X}(G_{p})$.
Case 2. Let $A_{r}={\sf Z}_{r^{k}}\times{\sf Z}_{r}$, where $k\geqslant 1$. Then $r$ does not divide $|B|$. By Lemma 4.3 (2), it implies that either
  • (i)
    $A_{r}={\sf Z}_{r}^{2}$ and $B$ is irreducible on $A_{r}$, or
  • (ii)
    there exists $a,b\in A_{r}$ such that $A_{r}=\langle a\rangle\times\langle b\rangle$, and both $\langle a\rangle$ and $\langle b\rangle$ are normalized by $B$.
For the former, $A_{r}$ centralizes $G_{p}$ since $A_{r}{:}B$ induces an abelian automorphism group of $G_{p}$ by Lemma 4.3. For the latter, there exist elements $u,v\in B$ such that $u$ acts on $\langle a\rangle$ non-trivially, and $v$ acts on $\langle b\rangle$ non-trivially. Then similar to Case 1 we have $G_{p},a]=[G_{p},b]=1$, so that, $G_{p},A_{r}]=[G_{p},\langle a,b\rangle]=1$.
We have therefore confirmed the claim. Consequently we have
$G=G_{p}{:}(A{:}B)=(G_{p}\times A){:}B.$
Meanwhile, by Lemma 4.3, $G_{p}$ is one of the following groups:
$G_{p}={\sf Z}_{p^{\ell}}$,  ${\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}$  or  ${\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$.
Note that by assumption, $G_{p^{\prime}}=A{:}B$ acts by conjugation on $G_{p}$ non-trivially, which implies that $B$ acts on $G_{p}$ non-trivially.
Suppose $G_{p}={\sf Z}_{p^{\ell}}$. As $B$ acts on $G_{p}$ non-trivially, it follows from Lemma 5.3 (ii) that $G_{p}\cap{\bf Z}(G)=1$. Thus G=C:DG=C{:}D satisfies the theorem, with $C=G_{p}\times A$, $D=B$.
Suppose $G_{p}={\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}=\langle a\rangle{:}\langle b\rangle$. By Lemma 4.3 (3), $B$ acts on $\langle a\rangle$ non-trivially and acts on $\langle b\rangle$ trivially. Rewrite $G$ as
$G=(A\times(\langle a\rangle{:}\langle b\rangle)){:}B=(A\times\langle a\rangle){:}(B\times\langle b\rangle).$
Then by Lemma 5.3 (ii), G=C:DG=C{:}D satisfies the theorem, with $C=A\times\langle a\rangle$, $D=B\times\langle b\rangle$.
Suppose $G_{p}={\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$, where $\ell\geqslant 1$. By Lemma 4.3 (2), it implies that either (i) $G_{p}={\sf Z}_{p}^{2}$ and $B$ is irreducible on $G_{p}$; or (ii) there exists $a,b\in G_{p}$ such that $G_{p}=\langle a\rangle\times\langle b\rangle$, and both $\langle a\rangle$ and $\langle b\rangle$ are normalized by $B$. For the former, G=C:DG=C{:}D satisfies the theorem, with $C=G_{p}\times A$, $D=B$. For the latter, if $B$ acts non-trivially on both $\langle a\rangle$ and $\langle b\rangle$, then also let $C=G_{p}\times A$, $D=B$. Otherwise, $B$ acting non-trivially on $G_{p}$ implies $B$ acting non-trivially on some one of $\langle a\rangle$ and $\langle b\rangle$, say $\langle a\rangle$. Then similar to the last paragraph, let $C=A\times\langle a\rangle$, $D=B\times\langle b\rangle$. $\Box$
We next deal with solvable groups of even order.
Lemma 5.5.
Let $G$ be a solvable group satisfying Hypothesis 2.2. If $F:={\rm Fit}(G)={\bf O}_{2}(G)$, then either $G$ is a $2$-group, or $G$ is one of the following:
  • (1)
    $G={\sf Z}_{2}^{3}{:}{\sf Z}_{7}$ or ${\sf Z}_{2}^{3}{:}{\sf Z}_{7}{:}{\sf Z}_{3};$
  • (2)
    $G={\sf Z}_{2}^{2}{:}H$, ${\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}H)$, ${\rm Q}_{8}{:}H$ or ${\sf Z}_{4}\circ({\rm Q}_{8}{:}H)$, where $H\in\{{\sf Z}_{3},{\rm S}_{3}\}$, ${\sf Z}_{2}^{2}{:}H\cong{\rm A}_{4}$ or ${\rm S}_{4}$, and ${\rm Q}_{8}{:}H\cong{\rm SL}(2,3)$ or ${\rm GL}(2,3)$.
Proof.
Assume that $G$ is not a 2-group. Then $G/F\leqslant{\sf Out}(F)$ is not a 2-group either. Thus by Lemma 4.1, $F$ is one of the following 2-groups:
$F\in\{{\sf Z}_{2}^{2},{\sf Z}_{2}^{3},{\rm Q}_{8},{\sf Z}_{4}\circ{\rm Q}_{8}\}.$
Moreover, ${\bf O}_{2}(G/F)$ is trivial, namely, $G/F$ has no normal 2-subgroups. Then by checking ${\sf Out}(F)$, either $|G/F|$ is divisible by 7, and $F={\sf Z}_{2}^{3}$, $G/F={\sf Z}_{7}$ or ${\sf Z}_{7}{:}{\sf Z}_{3}$; or $G/F={\sf Z}_{3}$ or ${\rm S}_{3}$. It follows that $G$ is one of the groups listed in the lemma. In particular, if $G=F{:}{\rm S}_{3}$, then $G_{2}=F{:}{\sf Z}_{2}>F$, and respectively for $F={\sf Z}_{2}^{2}$, ${\sf Z}_{2}^{3}$, ${\rm Q}_{8}$ or ${\sf Z}_{4}\circ{\rm Q}_{8}$, it is easy to verify that $G_{2}$ has a subgroup $K$ of index $2$ with $K={\sf Z}_{4}$, ${\rm D}_{8}$, ${\sf Z}_{8}$ or ${\rm D}_{16}$. That is, $G=F{:}{\rm S}_{3}$ surely satisfies Hypothesis 2.2. $\Box$
Lemma 5.6.
Let $K$ be a $\{2,p\}$-group with $p$ odd prime satisfying Hypotheses 2.2. If $K\neq K_{p}{:}K_{2}$, then one of the following holds.
  1. (1)
    $p=7$, $K={\sf Z}_{2}^{3}{:}K_{7};$
  2. (2)
    $p=3$, $K=X{:}K_{3}$ with $X\in\{{\sf Z}_{2}^{2},\,{\sf Z}_{2}^{3},\,{\rm Q}_{8},\,{\sf Z}_{4}\circ{\rm Q}_{8}\};$
  3. (3)
    $p=3$, $K$ is one of the groups in Tables 1 - 2
Proof.
Suppose that $K_{2}$ is normal in $K$, that is, $K=K_{2}{:}K_{p}$. Then $K_{2}$ has an automorphism of order $p$. By Lemma 4.1, $K_{2}$ is one of the following groups:
${\sf Z}_{2}^{2},\ {\rm Q}_{8},\ {\sf Z}_{2}^{3}\ \mbox{and}\ {\sf Z}_{4}\circ{\rm Q}_{8},$
and either $p=3$, $K=K_{2}{:}K_{3}$; or $K_{2}={\sf Z}_{2}^{3}$, $p=7$ and $K={\sf Z}_{2}^{3}{:}K_{7}$, as in (1) - (2).
We then assume that neither $K_{p}$ nor $K_{2}$ is normal in $K$. Let
$F={\rm Fit}(K)=F_{p}\times F_{2}$, with $F_{p}<K_{p}$, $F_{2}<K_{2}$.
Consider the quotient group $\overline{K}=K/F_{p}$, whose Fitting group is a 2-group, namely ${\rm Fit}(\overline{K})={\bf O}_{2}(\overline{K})$. Thus, by Lemma 5.5, $\overline{K}={\bf O}_{2}(\overline{K}){:}H$, where either
(i) ${\bf O}_{2}(\overline{K})\in\{{\sf Z}_{2}^{2},{\sf Z}_{2}^{3},{\rm Q}_{8},{\sf Z}_{4}\circ{\rm Q}_{8}\}$, $H\in\{{\sf Z}_{3},{\rm S}_{3}\}$; or (ii) ${\bf O}_{2}(\overline{K})={\sf Z}_{2}^{3}$, $H={\sf Z}_{7}$ or ${\sf Z}_{3}$.
Further, if (ii) holds, $\overline{K}={\sf Z}_{2}^{3}{:}{\sf Z}_{p}$ with $p=3$ or $7$, then ${\sf Z}_{2}^{3}$ is a minimal normal subgroup of $\overline{K}$, and so $\overline{F_{2}}=F_{2}$ as a normal subgroup of $\overline{K}$ must be identity. Consequently, we have $F=F_{p}$, ${\sf Z}_{2}^{3}{:}{\sf Z}_{p}=K/F_{p}\lesssim{\sf Out}(F_{p})$, which does not hold for the $p$-groups appearing in Lemma 4.3.
Therefore, $\overline{K}={\bf O}_{2}(\overline{K}){:}H$ is given in (i). We then analyze all the possibilities for $F_{2}=\overline{F_{2}}$, which a normal 2-subgroup of $\overline{K}$.
Case 1. $F_{2}={\bf O}_{2}(\overline{K})$. In this case, we have $H={\rm S}_{3}$ since $F_{2}<K_{2}$. Then
$K=F_{3}.(F_{2}{:}{\rm S}_{3})=(F_{3}\times F_{2}).{\rm S}_{3}=F_{2}{:}(F_{3}.{\rm S}_{3})=F_{2}{:}(K_{3}{:}{\sf Z}_{2}).$
Note that $K_{3}{:}{\sf Z}_{2}/F_{3}\cong{\rm S}_{3}$. By checking the $3$-groups appearing in Lemma 4.3, we conclude that there are seven possibilities for the triple $T:=(K_{3},\,K_{3}{:}{\sf Z}_{2},\,F_{3})$ as follows (also note that ${\sf Aut}({\sf Z}_{3}\times{\sf Z}_{3})={\rm GL}_{2}(3)$ is solvable):
[…]
Consequently, Case 1 gives the groups listed in Table 1.
Table 1. Case 1: $F_{2}={\bf O}_{2}(\overline{K})$
case $F_{2}$ ${\sf Z}_{2}^{2}$ ${\sf Z}_{2}^{3}$ ${\rm Q}_{8}$ ${\sf Z}_{4}\circ{\rm Q}_{8}$
(1.1) ${\rm S}_{4}$ ${\sf Z}_{2}\times{\rm S}_{4}$ ${\rm GL}_{2}(3)$ ${\sf Z}_{4}\circ{\rm GL}_{2}(3)$
(1.2) ${\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}}$ ${\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}})$ ${\rm Q}_{8}{:}{\rm D}_{2.3^{\ell+1}}$ ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell+1}})$
(1.3) ${\sf Z}_{3^{\ell}}\times{\rm S}_{4}$ ${\sf Z}_{2.3^{\ell}}\times{\rm S}_{4}$ ${\sf Z}_{3^{\ell}}\times({\rm Q}_{8}{:}{\rm S}_{3})$ ${\sf Z}_{4.3^{\ell}}\circ({\rm Q}_{8}{:}{\rm S}_{3})$
(1.4) ${\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})$ ${\sf Z}_{6}\times({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})$ ${\sf Z}_{3}\times({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})$ ${\sf Z}_{12}\circ({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})$
(1.5) $({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2}$ $({\sf Z}_{2.3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2}$ $({\sf Z}_{3^{\ell}}\times({\rm Q}_{8}{:}{\rm S}_{3})){:}{\sf Z}_{2}$ $({\sf Z}_{4.3^{\ell}}\circ({\rm Q}_{8}{:}{\rm S}_{3})){:}{\sf Z}_{2}$
(1.6) $({\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}$ $({\sf Z}_{6}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}$ $({\sf Z}_{3}\times({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}$ $({\sf Z}_{12}\circ({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}$
(1.7) ${\sf Z}_{2}^{2}{:}({\rm D}_{2.3^{\ell}}{:}{\sf Z}_{3})$ ${\sf Z}_{2}\times{\sf Z}_{2}^{2}{:}({\rm D}_{2.3^{\ell}}{:}{\sf Z}_{3})$ ${\rm Q}_{8}{:}({\rm D}_{2.3^{\ell}}{:}{\sf Z}_{3})$ ${\sf Z}_{4}\circ({\rm Q}_{8}{:}({\rm D}_{2.3^{\ell}}{:}{\sf Z}_{3}))$
Case 2. $(F_{2},{\bf O}_{2}(\overline{K}))=({\sf Z}_{2}^{2},{\sf Z}_{2}^{3})$ or $({\rm Q}_{8},{\sf Z}_{4}\circ{\rm Q}_{8})$. Then
$K=F_{3}.({\bf O}_{2}(\overline{K}){:}H)=(F_{3}\times F_{2}).({\sf Z}_{2}\times H)=F_{2}{:}(F_{3}.({\sf Z}_{2}\times H)).$
Further, if $H={\sf Z}_{3}$, then $K/F_{2}=K_{3}{:}{\sf Z}_{2}$. Note that ${\sf Z}_{2}$ acts non-trivially on $K_{3}$, and meanwhile it acts trivially on $K_{3}/F_{3}=H$. Then by checking the $3$-groups appearing in Lemma 4.3, we conclude that there are three possibilities:
[…]
Otherwise, $H={\rm S}_{3}$, and so $K/F_{2}=K_{3}{:}({\sf Z}_{2}\times{\sf Z}_{2})$. By Lemma 4.3, ${\sf Z}_{2}\times{\sf Z}_{2}\leqslant{\sf Aut}(K_{3})$ implies $K_{3}={\sf Z}_{3^{\ell}}\times{\sf Z}_{3}$ and $K_{3}{:}({\sf Z}_{2}\times{\sf Z}_{2})={\rm D}_{2.3^{\ell}}\times{\rm D}_{6}$. We then get two possibilities:
[…]
Consequently, Case 2 gives the groups listed in Table 2.
Table 2. Case 2: $F_{2}<{\bf O}_{2}(\overline{K})$
case $(F_{2},{\bf O}_{2}(\overline{K}))$ $({\sf Z}_{2}^{2},{\sf Z}_{2}^{3})$ $({\rm Q}_{8},{\sf Z}_{4}\circ{\rm Q}_{8})$
(2.1) ${\rm D}_{2.3^{\ell}}\times{\rm A}_{4}$ $({\sf Z}_{3^{\ell}}{:}{\sf Z}_{4})\circ({\rm Q}_{8}{:}{\sf Z}_{3})$
(2.2) $({\sf Z}_{2}^{2}{:}{{\sf Z}_{3^{\ell}}})\times{\rm D}_{6}$ $({\rm Q}_{8}{:}{{\sf Z}_{3^{\ell}}})\circ({\sf Z}_{3}{:}{\sf Z}_{4})$
(2.3) $({\rm D}_{2.3^{\ell}}\times{\sf Z}_{2}^{2}){:}{\sf Z}_{3}$ $(({\sf Z}_{3^{\ell}}{:}{\sf Z}_{4})\circ{\rm Q}_{8}){:}{\sf Z}_{3}$
(2.4) ${\rm D}_{2.3^{\ell}}\times{\rm S}_{4}$ $({\sf Z}_{3^{\ell}}{:}{\sf Z}_{4})\circ({\rm Q}_{8}{:}{\rm S}_{3})$
(2.5) $({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\rm D}_{6}$ $({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})\circ({\sf Z}_{3}{:}{\sf Z}_{4})$
Case 3. Recall that $F_{2}$ is a normal 2-subgroup of ${\bf O}_{2}(\overline{K})$. Notice that ${\rm Q}_{8}{:}{\sf Z}_{3}$ or ${\rm Q}_{8}{:}{\rm S}_{3}$ only has the center ${\sf Z}_{2}$ and ${\rm Q}_{8}$ as normal $2$-subgroups. Then all the remaining cases are collected as follows :
[…]
For case (3.1), $(F_{2},{\bf O}_{2}(\overline{K}))=(1,{\rm Q}_{8})$, we have
$F=F_{3}\mbox{\ and\ }K/F_{3}={\rm Q}_{8}{:}H\lesssim{\sf Aut}(F_{3}).$
By lemma 4.3, it implies that $F_{3}={\sf Z}_{3}\times{\sf Z}_{3}$ and $K_{3}={\sf Z}_{9}{:}{\sf Z}_{3}$ is non-abelian. Let
$K_{3}=\langle a\rangle{:}\langle b\rangle={\sf Z}_{9}{:}{\sf Z}_{3}$, with $a^{b}=a^{4}$, $b^{a}=a^{6}b$,
where $F_{3}=\langle a^{3}\rangle\times\langle b\rangle$. Note that $K_{2}=({\rm Q}_{8}{:}H)_{2}={\rm Q}_{8}$ or ${\rm Q}_{8}{:}{\sf Z}_{2}$. Consider the central involution $c$ of $K_{2}$. Then $\overline{c}$ is the central involution of $\overline{K}$. On the one hand, $\overline{c}$ acts on $F_{3}$ by mapping each $x$ to $x^{-1}$. That is,
$x^{c}=x^{-1}$, so that, $c^{x}=x^{-2}c$ for any $x\in F_{3}$.
On the other hand, $\overline{c}^{\overline{a}}=\overline{c}$, and so $c^{a}=cy$ where $y\in F_{3}$. Let $y=a^{3i}b^{j}$. Then
$c^{a^{3}}=(ca^{3i}b^{j})^{a^{2}}=(ca^{(6i+6j)}b^{2j})^{a}=ca^{(9i+18j)}b^{3j}=c,\mbox{\ where\ }a^{3}\in F_{3},$
a contradiction. Hence case (3.1) cannot occur.
For case (3.2), $(F_{2},{\bf O}_{2}(\overline{K}))=(1,{\sf Z}_{4}\circ{\rm Q}_{8})$, we have ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\sf Z}_{3})\lesssim{\sf Aut}(F_{3})$, which is impossible by checking the 3-groups appearing in Lemma 4.3.
For case (3.3), since ${\bf O}_{2}(\overline{K})$ acts on $F_{3}$ with the centralizer ${\bf C}_{{\bf O}_{2}(\overline{K})}(F_{3})=F_{2}$, we deduce that in this case ${\sf Aut}(F_{3})$ contains a subgroup isomorphic to ${\rm A}_{4}\cong{\sf Z}_{2}^{2}{:}{\sf Z}_{3}$. By Lemma 4.3, $|{\sf Aut}(F_{3})|$ being divisible by $4$ implies $F_{3}={\sf Z}_{3^{\ell}}\times{\sf Z}_{3}$. For $\ell=1$, it is clear that ${\sf Aut}({\sf Z}_{3}^{2})={\rm GL}(3,2)\cong{\rm Q}_{8}{:}{\rm S}_{3}$ contains no subgroups isomorphic to ${\rm A}_{4}$. We thus assume $\ell>1$ and let $F_{3}=\langle a\rangle\times\langle b\rangle$ with $|a|=3^{\ell}$, $|b|=3$. By Lemma 4.3, ${\sf Aut}(F_{3})$ is solvable, and it contains a Hall $3^{\prime}$-subgroup
$H_{0}=({\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle))_{3^{\prime}}\cong{\sf Z}_{2}^{2},$
which stabilizes subgroups $\langle a\rangle$ and $\langle b\rangle$. Meanwhile, by simple calculation, $H_{0}$ cannot stabilize
    any of the other two subgroups of order $3^{\ell}$, i.e., $\langle ab\rangle$, $\langle ab^{-1}\rangle$, and
    any of the other two non-characteristic subgroups of $3$, i.e., $\langle a^{3^{\ell-1}}b\rangle$, $\langle a^{-3^{\ell-1}}b\rangle$.
Consider the normalizer $N={\bf N}_{{\sf Aut}(F_{3})}(H_{0})$. For any $g\in N$, $H_{0}=H_{0}^{g}$ stabilizes subgroups $\langle a\rangle^{g}$ and $\langle b\rangle^{g}$, and so $\langle a\rangle^{g}=\langle a\rangle$, $\langle b\rangle^{g}=\langle b\rangle$, which implies that $N\leqslant{\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle)$ is abelian. Note that any subgroup $H\leqslant{\sf Aut}(F_{3})$ with $H\cong{\sf Z}_{2}^{2}$ is a Hall $3^{\prime}$-subgroup, then conjugate to $H_{0}$. Then the normalizer ${\bf N}_{{\sf Aut}(F_{3})}(H)$, which is conjugate to $N$, is also abelian. Thus, ${\sf Aut}(F_{3})$ has no subgroup isomorphic to ${\rm A}_{4}$.
Finally, with Case 3 excluded, the proof is complete. $\Box$
Lemma 5.7.
Let $K$ be a $\{2,3,7\}$-group satisfying Hypotheses 2.2, with neither $K_{3}$ nor $K_{7}$ normal in $K$. Then $K=K_{2}{:}(K_{7}{:}K_{3})={\sf Z}_{2}^{3}{:}(K_{7}{:}K_{3})$.
Proof.
By Lemma 5.1, $K_{\{3,7\}}=K_{7}{:}K_{3}$. Thus $K_{\{2,7\}}\neq K_{7}{:}K_{2}$, and by Lemma 5.6,
$K_{\{2,7\}}=K_{2}{:}K_{7}={\sf Z}_{2}^{3}{:}K_{7},$
Let $F={\rm Fit}(K)=F_{2}\times F_{3}\times F_{7}$. As $K_{2}$ is normalized by $K_{7}$, and meanwhile it normalizes $F_{7}$, so $K_{2}$ centralizes $F_{7}$. Moreover, $K_{2}$ also centralizes $F_{3}$. If not, consider the subgroup $F_{3}{:}(K_{2}{:}K_{7})\leqslant K$. Then ${\sf Z}_{2}^{3}$ as a minimal normal subgroup of $K_{2}{:}K_{7}$ is embedded into ${\sf Aut}(F_{3})$, which is impossible due to Lemma 4.3 and the fact ${\sf Z}_{2}^{3}\nleqslant{\rm GL}(2,3)\cong{\rm Q}_{8}{:}{\rm S}_{3}$. Hence $K_{2}$ centralizes $F$, so $K_{2}\leqslant C_{K}(F)\leqslant F$, and then $K_{2}=F_{2}\lhd K$, and $K=K_{2}{:}K_{\{3,7\}}=K_{2}{:}(K_{7}{:}K_{3})$. $\Box$
Now we are ready to prove Theorem 1.1.
Proof of Theorem 1.1:
Let $G$ be a finite solvable group of which each Sylow subgroup has a cyclic or dihedral subgroup of prime index. Let $H$ be the largest normal Hall subgroup of odd order, with $\pi=\pi(H)$. Let $K$ be a Hall $\pi^{\prime}$-subgroup of $G$, where $\pi^{\prime}=\pi(G)\setminus\pi(H)$. Then we have factorization
G=H:K.G=H{:}K.
Since $H$ is of odd order, by Lemma 5.4, H=A:BH=A{:}B such that $A$ is abelian, $B$ is nilpotent, and $A\cap{\bf Z}(H)=1$.
We then consider the subgroup $K$. Let $p$ be the largest prime divisor of $|G|$. If $p>7$, then by Lemmas 5.1 and 5.6, $G_{p}$ is normal in both $G_{2^{\prime}}$ and $G_{\{2,p\}}$, and so $G_{p}$ is normal in $G$, and $p\in\pi$. Then, by inductively considering $G/G_{p}$ we deduce that any prime divisor $r$ of $|G|$ with $r>7$ lies in $\pi$. That is, $\pi^{\prime}\subseteq\{2,3,5,7\}$. Further, by Lemmas 5.1, 5.2 and 5.6, $G_{5}$ is normalized by $G_{3}$, $G_{7}$ and $G_{2}$, respectively. Hence $5\notin\pi^{\prime}$. Thus, either
  • (i)
    $\pi^{\prime}=\emptyset$, $G=H$ is of odd order; or
  • (ii)
    $\pi^{\prime}=\{2\}$, $\{2,3\}$, $\{2,7\}$ or $\{2,3,7\}$.
At last, respectively for $\pi^{\prime}=\{2\}$, $\pi^{\prime}=\{2,3\}$ or $\{2,7\}$, and $\pi^{\prime}=\{2,3,7\}$, the group $K=G_{\pi^{\prime}}$ is described in Proposition 4.1, Lemma 5.6 and Lemma 5.7. $\Box$
6. Proof of Theorem 1.2
In this section we focus on the arc-transitive automorphism groups of maps. The following lemma shows that the class of such groups is closed under taking quotients.
Lemma 6.1.
Let $G$ be a finite group, with $\overline{G}$ a non-identity factor group. Then:
  • (1)
    If $G$ has a regular triple (resp. reversing triple), then either $\overline{G}$ has a regular triple (resp. reversing triple), or $\overline{G}$ is ${\sf Z}_{2}$ or a dihedral group. In particular, $\overline{G}$ is of even order.
  • (2)
    If $G$ has a rotary pair, then either $\overline{G}$ has a rotary pair, or $\overline{G}$ is cyclic. In particular, if $\overline{G}$ is of odd order, then it is cyclic.
Proof.
Let $(x,y,z)$ be a regular triple for $G$. If $\overline{x},\overline{y},\overline{z}\in\overline{G}$ are all involutions, and $\overline{x}\neq\overline{z}$, then they form a regular triple for $\overline{G}$, since $\overline{G}=\langle\overline{x},\overline{y},\overline{z}\rangle$ and $\overline{x}\,\overline{z}=\overline{z}\,\overline{x}$. Otherwise, either some of $\overline{x},\overline{y},\overline{z}$ equals the identity-element in $\overline{G}$, or $\overline{x}=\overline{z}$. In both cases, $\overline{G}$ is generated by one or two involutions and then it is ${\sf Z}_{2}$ or dihedral.
It is similar to prove the remaining part of this lemma. $\Box$
Lemma 6.2.
Let $G={\rm Q}_{8}{:}{\rm S}_{3}\cong{\rm GL}(2,3)$. Then neither $G$ nor ${\sf Z}_{4}\circ G$ has a regular triple.
Proof.
Suppose that $G$ has a regular triple $(x,y,z)$. Since $xz=zx$, we have $\langle x,z\rangle\cong{\sf Z}_{2}^{2}$, and so $\{x,z,xz\}\cap{\rm Q}_{8}\not$$\emptyset$. That is, one of $x,z,xz$ has to be the central involution 1-1 of $G$. It follows that $\overline{G}=G/\langle-1\rangle=\langle\overline{x},\overline{y}\rangle$ or $\langle\overline{y},\overline{z}\rangle$ is a dihedral group, which is impossible as $\overline{G}={\sf Z}_{2}^{2}{:}{\rm S}_{3}$.
Suppose that $K={\sf Z}_{4}\circ({\rm Q}_{8}{:}{\rm S}_{3})$ has a regular triple $(x,y,z)$. Write $K$ as
$K=\langle u\rangle\circ({\rm Q}_{8}{:}(\langle\sigma\rangle{:}\langle\tau\rangle)),$
where $|u|=4$, ${\rm Q}_{8}=\pm\{1,i,j,k\}$ with $i^{2}=j^{2}=k^{2}=ijk=-1$, and $\sigma,\tau\in{\sf Aut}({\rm Q}_{8})$ such that
$\sigma:i\rightarrow j\rightarrow k\,(\rightarrow i),\mbox{\,\ and\ \,}\tau:i\leftrightarrow-j,j\leftrightarrow-i,k\leftrightarrow-k.$
Then the subgroup $\langle\sigma\rangle{:}\langle\tau\rangle\cong{\rm S}_{3}$ contains three involutions, i.e., $\tau_{k}=\tau$,
$\tau_{i}:j\leftrightarrow-k,k\leftrightarrow-j,i\leftrightarrow-i,\mbox{\ and\ }\tau_{j}:i\leftrightarrow-k,k\leftrightarrow-i,j\leftrightarrow-j.$
By calculation, we conclude that $K$ contains the following involutions:
$\{-1\}\cup\{\pm u\lambda\}\cup\{\pm\tau_{\lambda},\pm\lambda\tau_{\lambda}\},\mbox{\ where\ }\lambda\in\{i,j,k\}.$
We shall confirm the following claim:
Claim. The only involutions commutes with $u\lambda$ (or $u\lambda$) are $\pm u\lambda$ and 1-1.
In fact, for any $\lambda^{\prime}\in\{i,j,k\}$ with $\lambda\neq\lambda^{\prime}$, we have:
$\lambda^{\lambda^{\prime}}=\lambda^{\tau_{\lambda}}=\lambda^{\lambda\tau_{\lambda}}=-\lambda$, and $\lambda^{\tau_{\lambda^{\prime}}}=-\lambda^{\prime\prime},\ \lambda^{\lambda^{\prime}\tau_{\lambda^{\prime}}}=-\lambda^{\tau_{\lambda^{\prime}}}=\lambda^{\prime\prime}$,
where $\lambda^{\prime\prime}\in\{i,j,k\}\setminus\{\lambda,\lambda^{\prime}\}.$ As $u$ and 1-1 lie in the center of $K$, we then have:
$(\pm u\lambda)^{\pm u\lambda^{\prime}}=\pm u(\lambda^{\pm u\lambda^{\prime}})=\pm u(\lambda^{\lambda^{\prime}})=\mp u\lambda,$
$(\pm u\lambda)^{\pm\tau_{\lambda}}=\pm u(\lambda^{\pm\tau_{\lambda}})=\mp u\lambda,\ (\pm u\lambda)^{\pm\lambda\tau_{\lambda}}=\pm u(\lambda^{\pm\lambda\tau_{\lambda}})=\mp u\lambda,$
$(\pm u\lambda)^{\pm\tau_{\lambda^{\prime}}}=\pm u(\lambda^{\pm\tau_{\lambda^{\prime}}})=\mp u\lambda^{\prime\prime},\ (\pm u\lambda)^{\pm\lambda^{\prime}\tau_{\lambda^{\prime}}}=\pm u(\lambda^{\pm\lambda^{\prime}\tau_{\lambda^{\prime}}})=\pm u\lambda^{\prime\prime}.$
We have therefore confirmed the claim.
Let $H=\langle u,i,j,k\rangle\leqslant K$. Then $H\cong{\sf Z}_{4}\circ{\rm Q}_{8}$ is normal in $K$ and $K/H\cong{\rm S}_{3}$. Further, the involutions contained in $H$ are: $\{-1\}\cup\{\pm u\lambda\}$. Recall that $(x,y,z)$ is a regular triple for $G$. Thus, $\langle x,z\rangle\cong{\sf Z}_{2}^{2}$, and so $\langle x,z\rangle\cap H>1$, that is,
$\{x,z,xz\}\cap(\{-1\}\cup\{\pm u\lambda\})\not$$\emptyset$.
By the claim given above, one of $x,z,xz$ has to be 1-1. Thus, $\overline{K}=K/\langle-1\rangle=\langle\overline{x},\overline{y}\rangle$ or $\langle\overline{y},\overline{z}\rangle$ is a dihedral group, which is impossible as $\overline{K}=K/\langle-1\rangle={\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}{\rm S}_{3})$. $\Box$
Lemma 6.3.
Let $G=({\sf Z}_{3^{\ell}}\times{\sf Z}_{3}){:}{\sf Z}_{2}=\langle a,b\rangle{:}\langle c\rangle$, with $a^{c}=a^{-1}$, $b^{c}=b^{-1}$. Then $G$ has no rotary pair.
Proof.
Let $(\alpha,z)$ be a rotary pair for $G$. Since $G$ is not dihedral, $|\alpha|\neq 2$, so $\alpha\in\langle a,b\rangle$ and $\langle\alpha\rangle$ is normal in $G$. It follows that $\overline{G}=G/\langle\alpha\rangle=\langle\overline{z}\rangle={\sf Z}_{2}$, not possible. $\Box$
Now we are ready to prove Theorem 1.2.
Proof of Theorem 1.2:
Let $\mathcal{M}$} be a map with square-free Euler characteristic, and let $G\leqslant{\sf Aut}{\mathcal{M}}$ be a solvable arc-transitive automorphism group of $\mathcal{M}$}. By Lemmas 2.1 and 2.3, the group $G$ satisfies Hypothesis 2.2, so that
G=(A:B):KG=(A{:}B){:}K
as described in Theorem 1.1.
Recall that $\mathcal{M}$} as a $G$-arc-transitive map satisfies one of the following properties: (i) $G$-regular; (ii) $G$-vertex-reversing; (iii) $G$-vertex-rotary; respectively in each case, the group $G$ has: (i) a regular triple; (ii) a reversing triple; (iii) a rotary pair. In particular, $G$ is of even order, and so $K\neq 1$.
First, suppose $K=G_{2}$. By Lemma 6.1, either $K$ has a regular triple/reversing triple/rotary pair as $G$ does, and then $K$ is one of the 2-groups given in Proposition 4.2; or $K$ is ${\sf Z}_{2}$, dihedral ${\rm D}_{2^{\ell}}$ or cyclic ${\sf Z}_{2^{\ell}}$ with $\ell\geqslant 2$, while ${\rm D}_{2^{\ell}}$ and ${\sf Z}_{2^{\ell}}$ already appear in Proposition 4.2.
Then, suppose $K=G_{\{2,7\}}={\sf Z}_{2}^{3}{:}G_{7}$ or $K=G_{\{2,3,7\}}={\sf Z}_{2}^{3}{:}(G_{7}{:}G_{3})$. In this case, $G$ has an odd factor group $G_{7}$ or $G_{7}{:}G_{3}$, respectively. By Lemma 6.1, G has neither regular triples nor reversing triples, so $G$ has a rotary pair, with each odd factor group cyclic. It follows that $K=G_{\{2,7\}}={\sf Z}_{2}^{3}{:}{\sf Z}_{7^{\ell}}$.
At last, suppose that $K=G_{\{2,3\}}$ is one of the groups given in 5.6 (2) - (3).
(a) If $\mathcal{M}$} is $G$-vertex-reversing, then $G$ has a reversing triple, with each factor group of even order by Lemma 6.1. Hence, $K$ is one of the groups in Tables 1 - 2, excluding those corresponding to cases (1.3), (1.4), (1.7), (2.1) - (2.3) as they have an odd factor group ${\sf Z}_{3}$ or ${\sf Z}_{3^{\ell}}$.
(b) If $\mathcal{M}$} is $G$-regular, then $G$ has a regular triple. In particular, $G$ has a reversing triple, and so $K$ is one of the groups in (a). Furthermore, by Lemma 6.2 the groups which have ${\rm Q}_{8}{:}{\rm S}_{3}$ or ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\rm S}_{3})$ as a factor group should be excluded. It follows that $K$ is one of: $X$, $X\times{\sf Z}_{2}$, ${\rm D}_{2.3^{\ell}}\times{\rm S}_{4}$ or $({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\rm D}_{6}$, where
$X\in\{{\rm S}_{4},\,{\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}},\,({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2},\,({\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}\}.$
(c) If $\mathcal{M}$} is $G$-vertex-rotary, then $G$ has a rotary pair, with each odd factor group cyclic by Lemma 6.1. So, either $K=X{:}K_{3}=X{:}{\sf Z}_{3^{\ell}}$ with $X\in\{{\sf Z}_{2}^{2},\,{\sf Z}_{2}^{3},\,{\rm Q}_{8},\,{\sf Z}_{4}\circ{\rm Q}_{8}\};$ or $K$ is one of the groups in Tables 1 - 2. For the former, $X\neq{\rm Q}_{8}$, as the only involution of ${\rm Q}_{8}{:}{\sf Z}_{3^{\ell}}$ lies in the center and so it does not have a rotary pair. For the latter, by Lemma 6.3 the groups corresponding to cases (1.5) and (1.6), which have $({\sf Z}_{3^{\ell}}\times{\sf Z}_{3}){:}{\sf Z}_{2}$ as a factor group, should be excluded.
In conclusion, $K$ is one of the groups stated in Theorem 1.2. $\Box$
Declarations
Funding and/or Conflicts of interests/Competing interests
The author(s) declares that there is no any financial or personal relationship with other people or organizations not mentioned that can inappropriately influence the work.